### v2

# R-matrix valued Lax pair for elliptic Calogero-Inozemtsev system

Abstract
We consider the elliptic Calogero-Inozemtsev system of ${\rm BC}_{n}$ type with five arbitrary constants and propose $R$-matrix valued generalization for $2n\times 2n$ Takasaki’s Lax pair. For this purpose we extend the Kirillov’s ${\rm B}$-type associative Yang-Baxter equations to the similar relations depending on the spectral parameters and the Planck constants. General construction uses the elliptic Shibukawa-Ueno $R$-operator and the Komori-Hikami $K$-operators satisfying reflection equation. Then, using the Felder-Pasquier construction the answer for the Lax pair is also written in terms of the Baxter’s 8-vertex $R$-matrix. As a by-product of the constructed Lax pair we also propose ${\rm BC}_{n}$ type generalization for the elliptic XYZ long-range spin chain, and we present arguments pointing to its integrability.
1 Introduction
Calogero-Moser model and $R$-matrices.
The elliptic Calogero-Moser model of ${\rm gl}_{n}$ type is an integrable systems of classical mechanics, describing pairwise interaction of $n$ particles. It is defined by the Hamiltonian function [25]:
$\begin{array}[]{c}\displaystyle{H=\sum\limits_{i=1}^{n}\frac{p_{i}^{2}}{2}-{g}^{2}\sum\limits_{i>j}^{n}\wp(q_{i}-q_{j})\,,}\end{array}$ (1.1)
where $\wp(z)$ is the Weierstrass Weierstrass-$\wp$-function, $g\in\mathbb{C}$ is a coupling constant, $p_{i}\in\mathbb{C}$ and $q_{i}\in\mathbb{C}$ are the particles momenta and positions. The Poisson brackets are canonical:
$\begin{array}[]{c}\displaystyle{\{p_{i},q_{j}\}=\delta_{ij}\,,\quad\{p_{i},p_{j}\}=\{q_{i},q_{j}\}=0\,.}\end{array}$ (1.2)
The equations of motion
$\begin{array}[]{c}\displaystyle{{\dot{q}}_{i}=p_{i}\,,\quad{\ddot{q}}_{i}={g}^{2}\sum\limits_{k:k\neq i}^{n}\wp^{\prime}(q_{ik})}\end{array}$ (1.3)
are represented in the Lax form
$\begin{array}[]{c}\displaystyle{{\dot{L}}(z)\equiv\{H,L(z)\}=[L(z),M(z)]}\end{array}$ (1.4)
with spectral parameter $z\in\mathbb{C}$ (a coordinate on elliptic curve with moduli $\tau$, ${\rm Im}(\tau)>0$). Explicit form for the Lax pair $L(z),M(z)$ was found in [19] by I. Krichever. The Lax matrix is of the size $n\times n$:
$\begin{array}[]{c}\displaystyle{L(z)=\sum\limits_{i,j=1}^{n}E_{ij}\,L_{ij}(z)\,,\quad L_{ij}(z)=\delta_{ij}p_{i}+{g}(1-\delta_{ij})\phi(z,q_{ij})\,,\quad q_{ij}=q_{i}-q_{j}\,,}\end{array}$ (1.5)
where $\phi(z,u)$ is the elliptic Kronecker function [32]
$\begin{array}[]{l}\displaystyle{\phi(z,u)=\frac{\vartheta^{\prime}(0)\vartheta(z+u)}{\vartheta(z)\vartheta(u)}}\end{array}$ (1.6)
defined through $\vartheta(z)$ – the first Jacobi theta-function (A.2). The expression for the $M$-matrix is given in the next Section. The validity of the Lax equation is based on a set of elliptic function identities for the function $\phi(z,u)$. The main one is the following summation formula (or the genus one Fay identity, see (A.12)):
$\begin{array}[]{c}\displaystyle{\phi(z,q_{ik})\phi(w,q_{kj})=\phi(w,q_{ij})\phi(z-w,q_{ik})+\phi(w-z,q_{jk})\phi(z,q_{ij})\,.}\end{array}$ (1.7)
This relation is a particular case of a more general one known as the associative Yang-Baxter equation (AYBE) [7, 27]:
$\begin{array}[]{c}\displaystyle{R^{z}_{12}R^{w}_{23}=R^{w}_{13}R_{12}^{z-w}+R^{w-z}_{23}R^{z}_{13}\,,\ \ R^{z}_{ij}=R^{z}_{ij}(q_{i},q_{j})\,,}\end{array}$ (1.8)
where $R^{z}_{12}(x_{1},x_{2})$ is an operator that acts on a product of vector spaces. For $R$-matrices depending on the difference of spectral parameters $R^{z}_{12}(x_{1},x_{2})=R^{z}_{12}(x_{1}-x_{2})$ and in the case where the action on vector spaces is by multiplication on the function $\phi(z,x_{1}-x_{2})$ we reproduce (1.7). In the general case, equation (1.8) is fulfilled by the elliptic Baxter-Belavin $R$-matrix [27].
It can be show (see e.g. [21]) that any solution of (1.8) satisfying certain additional properties (see below the unitarity (2.5) the and skew-symmetry (2.6)) satisfies also the quantum Yang-Baxter equation
$\begin{array}[]{l}\displaystyle{R^{\hbar}_{12}(x_{1},x_{2})R^{\hbar}_{13}(x_{1},x_{3})R^{\hbar}_{23}(x_{2},x_{3})=R^{\hbar}_{23}(x_{2},x_{3})R^{\hbar}_{13}(x_{1},x_{3})R^{\hbar}_{12}(x_{1},x_{2})\,.}\end{array}$ (1.9)
For this reason, we refer to any solution of the AYBE (1.8) as $R$-matrix [^1] In fact, the Yang-Baxter equations (1.8) and (1.9) have different (and intersecting) sets of solutions. We consider a special class of solutions of AYBE (1.8), which also satisfy the quantum Yang-Baxter equation (1.9).. Many possible applications of the AYBE to different algebraic and geometric constructions, including those related to integrable systems, can be found in [7, 27, 26, 18] and [21, 22, 23, 28, 9, 34, 24].
The similarity between (1.7) and (1.8) was used in [21] to propose the so-called $R$-matrix valued Lax pair, which is a Lax matrix with entries
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}_{ij}(z)=\delta_{ij}p_{i}+{g}(1-\delta_{ij})R^{z}_{ij}(q_{i}-q_{j})\,.}\end{array}$ (1.10)
Different applications and properties of this Lax representation can be found in [21, 22, 23, 28, 9, 34]. In particular, in [28] the anisotropic (XYZ) version of the Haldane-Shastry-Inozemtsev long-range spin chain was proposed using $R$-matrix valued Lax pair.
${\rm BC}_{n}$ type Calogero-Inozemtsev system and reflection equation.
The elliptic Calogero-Moser model (1.1) is related to the root system ${\rm A}_{n-1}$ and can be extended to other root systems [25, 4]. The model with five arbitrary constants was suggested by V. Inozemtsev [13]. It is described by the Hamiltonian:
${H=\frac{1}{2}\sum_{k=1}^{n}p_{k}^{2}-g^{2}\sum_{i<j}^{n}\Big{(}\wp(q_{i}-q_{j})+\wp(q_{i}+q_{j})\Big{)}-\frac{1}{2}\sum\limits_{a=0}^{3}\sum_{k=1}^{n}\nu_{a}^{2}\wp(q_{k}+\omega_{a})\,,}$ (1.11)
where $\omega_{\gamma}$ are half-periods (A.18), and the five arbitrary constants are $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$. Originally, the Lax representation was of size $3n\times 3n$. We use $2n\times 2n$ Lax representation proposed by K. Takasaki in [31]. More precisely, we use it in the form obtained by O. Chalykh in [3] as a limiting case of a more general relativistic (Ruijsenaars-van Diejen) model.
In our construction of the $R$-matrix valued Lax pair we also use the $K$-matrices [12], which are solutions to the reflection equation [30]:
$\begin{array}[]{c}\displaystyle{R^{-}_{12}(x_{1},x_{2})K^{\hbar}_{1}(x_{1})R^{+}_{12}(x_{1},x_{2})K^{\hbar}_{2}(x_{2})=K^{\hbar}_{2}(x_{2})R^{+}_{12}(x_{1},x_{2})K^{\hbar}_{1}(x_{1})R^{-}_{12}(x_{1},x_{2})\,,}\end{array}$ (1.12)
where $R^{-}_{12}(x_{1},x_{2})=R^{\hbar}_{12}(x_{1}-x_{2})$, $R^{+}_{12}(x_{1},x_{2})=R^{\hbar}_{12}(x_{1}+x_{2})$. The $K$-matrices play the role of the boundary conditions in quantum integrable systems.
In [11, 16, 17] Y. Komori and K. Hikami suggested solution of (1.12) with the operator valued Shibukawa-Ueno $R$-matrix in the form of certain operators ($K$-operators). This was applied for constructing commuting set of quantum Hamiltonians for the model (1.11) and its relativistic generalization (through the Dunkl-Cherednik type operators). We actively use the results of [11, 16, 17] in our consideration.
Purpose of the paper.
The associative Yang-Baxter equation was written in the form (1.8) in [27], and is also known as the associative Yang-Baxter equation with parameters, while the original equation [7] appeared in the form of a set of relations (details are given in Section 4) including the main one
$\begin{array}[]{c}\displaystyle{r_{ij}r_{jk}=r_{ik}r_{ij}+r_{jk}r_{ik}\qquad\text{ for distinct }i,j,k\,.}\end{array}$ (1.13)
Then (1.8) can be considered as representation of (1.13) with parameters. In [18] the set of relations was extended to other root systems. For the root system of type $B$ the set of relations includes the four term relation
${r_{ij}y_{j}=y_{i}r_{ij}+\tilde{r}_{ij}y_{i}+y_{j}\tilde{r}_{ij}\,.}$ (1.14)
In Section 4 we propose its analogue with parameters and call it the ${BC}_{n}$ associative Yang-Baxter equation with parameters. In particular, the analogue of the four term relation (1.14) is as follows:
${R_{ij}^{w+z}(q_{i}-q_{j})\tilde{K}^{w}_{j}(q_{j})=\tilde{K}^{w}_{i}(q_{i})R_{ij}^{z-w}(q_{i}-q_{j})+\tilde{R}_{ij}^{w-z}(q_{i}+q_{j})\tilde{K}^{z}_{i}(q_{i})+\tilde{K}_{j}^{-z}(q_{j})\tilde{R}_{ij}^{w+z}(q_{i}+q_{j})\,.}$ (1.15)
To obtain this result we use the operator-valued Shibukawa-Ueno $R$-matrix and the construction similar to the Komori-Hikami $K$-matrix.
Next, we show that the obtained relations allow to construct $R$-matrix valued Lax pair for the Calogero-Inozemtsev model by generalizing the Takasaki’s $2n\times 2n$ Lax representation with spectral parameter.
It was explained by G. Felder and V. Pasquier in [6] that the Shibukawa-Ueno $R$-operator can be transformed into the elliptic Baxter-Belavin $R$-matrix by specifying the space, where this operator acts. In a similar way the $K$-operators from [11, 16] were transformed into the matrix form in [17]. We use these results and prove that the obtained relations for the ${\rm BC}_{n}$ associative Yang-Baxter equation with parameters are valid for the Baxter’s 8-vertex $R$-matrix and the elliptic $K$-matrix.
Finally, using the approach of [28] we present the quantum Hamiltonian for anisotropic long-range spin chain of ${\rm BC}_{n}$ type.
2 Elliptic Calogero-Moser model and $R$-matrices
2.1 Krichever’s Lax pair
Consider the Lax pair for the elliptic ${\rm gl}_{n}$ Calogero-Moser model (1.1)-(1.3). The Lax pair of size $n\times n$, and the Lax matrix has the form (1.5) Similarly, for the matrix elements of $M$-matrix:
$\begin{array}[]{c}\displaystyle{M_{ij}(z)={g}d_{i}\delta_{ij}+{g}(1-\delta_{ij})f(z,q_{ij})\,,\quad d_{i}=\sum\limits_{k:k\neq i}^{n}E_{2}(q_{ik})=-\sum\limits_{k:k\neq i}^{n}f(0,q_{ik})\,,}\end{array}$ (2.1)
See Appendix for elliptic functions definitions and properties. The proof of the Lax equation (1.4) with the Lax pair (1.5), (2.1) is based on the identities (A.15)-(A.17) written as
$\begin{array}[]{c}\displaystyle{\phi(z,q_{ik})f(z,q_{kj})-f(z,q_{ik})\phi(z,q_{kj})=\phi(z,q_{ij})(f(0,q_{kj})-f(0,q_{ik}))\,.}\end{array}$ (2.2)
and
$\begin{array}[]{c}\displaystyle{\phi(z,q_{ij})f(z,q_{ji})-f(z,q_{ij})\phi(z,q_{ji})=\wp^{\prime}(q_{ij})\,.}\end{array}$ (2.3)
These relations follow from the (genus one Fay identity) summation formula (A.12):
$\begin{array}[]{c}\displaystyle{\phi(z,q_{ik})\phi(w,q_{kj})=\phi(w,q_{ij})\phi(z-w,q_{ik})+\phi(w-z,q_{jk})\phi(z,q_{ij})\,.}\end{array}$ (2.4)
2.2 $R$-matrix identities
The elliptic $R$-matrix satisfies the unitarity property
$\begin{array}[]{l}\displaystyle{R_{12}^{\hbar}(z)R_{21}^{\hbar}(-z)=(\wp(\hbar)-\wp(z))\,.}\end{array}$ (2.5)
This is an $R$-matrix analogue of (A.16). Also, we have a skew-symmetry property
${R^{\hbar}_{12}(z)=-R^{-\hbar}_{21}(-z)\,,}$ (2.6)
which is an analogue of the simple relation $\phi(\hbar,z)=-\phi(-\hbar,-z)$. Main relation is the associative Yang-Baxter equation [7]
$\begin{array}[]{c}\displaystyle{R^{z}_{ik}R^{w}_{kj}=R^{w}_{ij}R_{ik}^{z-w}+R^{w-z}_{kj}R^{z}_{ij}\,,\ \ R^{z}_{ij}=R^{z}_{ij}(q_{i}\!-\!q_{j})\,.}\end{array}$ (2.7)
This is a matrix generalization of the Fay identity (A.12) or (2.4), and it is fulfilled by the elliptic Baxter-Belavin $R$-matrix [27]. Introduce also matrix analogue of the function $f(z,u)$ (A.7):
$\begin{array}[]{c}\displaystyle{F^{\,z}_{ij}(u)=\partial_{u}R^{\,z}_{ij}(u)}\end{array}$ (2.8)
and
[…]
Then similarly to (A.15) or (2.2) one gets
$\begin{array}[]{c}\displaystyle{R^{z}_{ik}F^{z}_{kj}-F^{z}_{ik}R^{z}_{kj}=F^{0}_{kj}R^{z}_{ij}-R^{z}_{ij}F^{0}_{ik}\,.}\end{array}$ (2.10)
By differentiating (2.5) we also obtain
$\begin{array}[]{c}\displaystyle{R^{z}_{ij}F^{z}_{ji}-F^{z}_{ij}R^{z}_{ji}=\wp^{\prime}(q_{ij})\,.}\end{array}$ (2.11)
2.3 $R$-matrix valued Lax pairs
The Lax pair (1.5), (2.1) of the ${\rm gl}_{n}$ Calogero-Moser model has the following generalization [21] (see also [10, 28, 34]) called the $R$-matrix valued Lax pair:
$\begin{array}[]{c}\displaystyle{{{\mathcal{L}}}(z)=\sum\limits_{i,j=1}^{n}E_{ij}\otimes{\mathcal{L}}_{ij}(z)\,,\quad\quad{\mathcal{L}}_{ij}(z)=1_{N}^{\otimes n}\,\delta_{ij}p_{i}+g(1-\delta_{ij})R^{\,z}_{ij}(q_{ij})}\end{array}$ (2.12)
¯$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}_{ij}(z)=D_{i}\delta_{ij}+g(1-\delta_{ij})F^{\,z}_{ij}(q_{ij})+\delta_{ij}\,{\mathcal{F}}^{\,0}\,,\quad D_{i}=-g\sum\limits_{k:k\neq i}^{n}F^{\,0}_{ik}(q_{ik})\,,}\end{array}$ (2.13)
where
[…]
The proof is based on the identities (2.10)-(2.11) and the following simple [^2] Relation (2.15) does not use any specific $R$-matrix properties but only $[R_{ab}^{z},F^{w}_{cd}]=0$ for all distinct $a,b,c,d$. relation:
$\begin{array}[]{c}\displaystyle{[R^{z}_{ij},{\mathcal{F}}^{\,0}]+\sum\limits_{l:l\neq i,j}R^{z}_{il}F^{z}_{lj}-F^{z}_{il}R^{z}_{lj}=\sum\limits_{l:l\neq j}R^{z}_{ij}F^{0}_{lj}-\sum\limits_{l:l\neq i}F^{0}_{il}R^{z}_{ij}=D_{i}R_{ij}^{z}-R_{ij}^{z}D_{j}\,,\quad\forall\ i\neq j.}\end{array}$ (2.15)
Similarly, $R$-matrix valued Lax pairs can be deduced for other root systems. Such results were described in [10] by extending the Lax pairs [25, 4] for the Calogero-Moser models related to the classical root systems of $B,C,D$ types. In accordance with the original construction from [25, 4] in these cases we have not arbitrary coupling constants, they satisfy a set of constraints. In [22] the ${BC}_{1}$ case was considered with four arbitrary constants. In both papers [10] and [22] an additional component of the quantum Hilbert space was used in order to define matrix elements ${\mathcal{L}}_{i,i+n}$ and ${\mathcal{L}}_{n+i,i}$. Below we use different approach based on the usage of $K$-matrices.
2.4 Long-range spin chains
Having an $R$-matrix valued Lax pair one can deduce a quantum long-range spin chain in the following way [28]. Denote by ${{\mathcal{M}}}(z)$ the part of ¯$\bar{\mathcal{M}}(z)$ without the term ${\mathcal{F}}^{0}$, that is
¯$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}(z)={\mathcal{M}}(z)+1_{n}{\mathcal{F}}^{0}\,.}\end{array}$ (2.16)
Next, consider the Lax equation
˙$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),\bar{\mathcal{M}}(z)]}\end{array}$ (2.17)
at an equilibrium position, where $p_{i}={\dot{q}}_{i}=0$. The positions of particles are fixed at some points satisfying equations ${\dot{p}}_{i}=0$. For the elliptic Calogero-Moser model (1.1) this set of points is
$\begin{array}[]{c}\displaystyle{q_{j}\rightarrow x_{j}=\frac{j}{n}\,,\quad j=1,...,n\,.}\end{array}$ (2.18)
Consider restrictions of the Lax matrices to the equilibrium position:
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{q_{j}=x_{j},p_{j}=0}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=x_{j}}\,,\quad{{\mathcal{F}}^{0}}^{\prime}={\mathcal{F}}^{0}|_{q_{j}=x_{j}}\,.}\end{array}$ (2.19)
Then from the Lax equation (taking into account that ˙${\dot{\mathcal{L}}}(z)|_{q_{j}=x_{j}}=0$) we get
$\begin{array}[]{c}\displaystyle{[{{\mathcal{F}}^{0}}^{\prime},{\mathcal{L}}^{\prime}(z)]=[{\mathcal{L}}^{\prime}(z),{\mathcal{M}}^{\prime}(z)]\,,}\end{array}$ (2.20)
which is a quantum Lax equation for a quantum model defined by the Hamiltonian ${{\mathcal{F}}^{0}}^{\prime}$.
This construction can be regarded as an explanation of the Polychronakos freezing trick, which relates the quantum spin Calogero-Moser models with the long-range spin chains by restricting to the classical equilibrium position of the corresponding spinless model. The above idea was used in [28] to obtain anisotropic version of the Inozemtsev long-range chain [14]. Let us also mention that isotropic trigonometric ${\rm BC}_{n}$ type long-range spin chains were described in [5] using the Dunkl operator formalism.
Different generalizations and applications for this type of models can be found in [20, 24, 15] and references therein.
3 Takasaki’s $2n\times 2n$ Lax pair for Calogero-Inozemtsev system
3.1 The model and the Lax pair
Here we recall the Lax pair for the Calogero-Inozemtsev model (1.11) with five arbitrary constants $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$ Equations of motion are of the form:
${{\dot{p}}_{i}={\ddot{q}}_{i}=g^{2}\sum_{k:k\neq i}^{n}\Big{(}\wp^{\prime}(q_{i}-q_{k})+\wp^{\prime}(q_{i}+q_{k})\Big{)}+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(q_{i}+\omega_{a})\,.}$ (3.1)
In [13] $3n\times 3n$ Lax representation was suggested. Some particular cases (not arbitrary constants) were known previously from [25], see also [4]. Here we are going to deal with $2n\times 2n$ Lax representation proposed by K. Takasaki in [31]. See also [3], where this Lax pair arises in the non-relativistic limit, and we use the expression from [3]. The description of the spectral curve can be found in [1], and the case of $n=1$ was also studied in [33].
The Lax pair has a natural block-matrix structure:
$\begin{array}[]{c}\displaystyle{L(z)=\left(\begin{array}[]{cc}{L^{11}(z)}&{L^{12}(z)}\\ \\ {L^{21}(z)}&{L^{22}(z)}\end{array}\right)\,,\qquad M(z)=\left(\begin{array}[]{cc}{M^{11}(z)}&{M^{12}(z)}\\ \\ {M^{21}(z)}&{M^{22}(z)}\end{array}\right)\,,}\end{array}$ (3.2)
where all entries $L^{ab}(z)$ and $M^{ab}(z)$, $a,b=1,2$ are $n\times n$ matrices (below $i,j=1,...,n$):
$\begin{array}[]{c}\displaystyle{L^{11}_{ij}(z)=\delta_{ij}p_{i}+g(1-\delta_{ij})\phi(z,q_{ij})\,,}\end{array}$ (3.3)
$\begin{array}[]{c}\displaystyle{L^{12}_{ij}(z)=\delta_{ij}v(z,q_{i})+g(1-\delta_{ij})\phi(z,q^{+}_{ij})\,,}\end{array}$ (3.4)
$\begin{array}[]{c}\displaystyle{L^{21}_{ij}(z)=-\delta_{ij}v(-z,q_{i})-g(1-\delta_{ij})\phi(-z,q^{+}_{ij})\,,}\end{array}$ (3.5)
$\begin{array}[]{c}\displaystyle{L^{22}_{ij}(z)=-\delta_{ij}p_{i}-g(1-\delta_{ij})\phi(-z,q_{ij})}\end{array}$ (3.6)
and
$\begin{array}[]{c}\displaystyle{M^{11}_{ij}(z)=\delta_{ij}a_{i}+g(1-\delta_{ij})f(z,q_{ij})\,,}\end{array}$ (3.7)
$\begin{array}[]{c}\displaystyle{M^{12}_{ij}(z)=\frac{1}{2}\,\delta_{ij}v^{\prime}(z,q_{i})+g(1-\delta_{ij})f(z,q^{+}_{ij})\,,}\end{array}$ (3.8)
$\begin{array}[]{c}\displaystyle{M^{21}_{ij}(z)=\frac{1}{2}\,\delta_{ij}v^{\prime}(-z,q_{i})+g(1-\delta_{ij})f(-z,q^{+}_{ij})\,,}\end{array}$ (3.9)
$\begin{array}[]{c}\displaystyle{M^{22}_{ij}(z)=\delta_{ij}a_{i}+g(1-\delta_{ij})f(-z,q_{ij})\,,}\end{array}$ (3.10)
where the functions $\phi$ and $f$ are given by (2.5) and (A.7) respectively and the following notations are used:
$\begin{array}[]{c}\displaystyle{q_{ij}=q_{i}-q_{j}\,,\quad q^{+}_{ij}=q_{i}+q_{j}\,.}\end{array}$ (3.11)
The block-matrix (3.2) has the following (anti)symmetry properties:
$L^{11}(z)=-L^{22}(-z),\qquad L^{12}(z)=-L^{21}(-z)\,,$ (3.12)
$M^{11}(z)=M^{22}(-z),\qquad M^{12}(z)=M^{21}(-z)\,.$ (3.13)
The diagonal elements of $M^{11}(z)$ and $M^{22}(z)$ are given by expressions
$\begin{array}[]{c}\displaystyle{a_{i}=g\sum\limits_{k:k\neq i}^{n}\Big{(}\wp(q_{i}-q_{k})+\wp(q_{i}+q_{k})\Big{)}+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}\wp(q_{i}+\omega_{a})}\end{array}$ (3.14)
with half-periods $\omega_{a}$ numerated as in (A.18). Using the set of functions (A.18)-(A.19) we also define
$\begin{array}[]{c}\displaystyle{v(z,u)\equiv v(z,u|\nu)=\sum\limits_{a=0}^{3}\nu_{a}\varphi_{a}(2z,u+\omega_{a})=\sum\limits_{a=0}^{3}\nu_{a}\exp(4\pi\imath z\partial_{\tau}\omega_{a})\phi(2z,u+\omega_{a})}\end{array}$ (3.15)
and its partial derivative
$\begin{array}[]{c}\displaystyle{v^{\prime}(z,u)=\partial_{u}v(z,u)=\sum\limits_{a=0}^{3}\nu_{a}\exp(4\pi\imath z\partial_{\tau}\omega_{a})f(2z,u+\omega_{a})\,.}\end{array}$ (3.16)
Some useful formulae for the functions $v(z,u)$ and $v^{\prime}(z,u)$ can be found in [16, 17, 11] and [33]. In particular, we have
$\begin{array}[]{c}\displaystyle{v(z,u)v(z,-u)=\sum\limits_{a=0}^{3}\Big{(}\bar{\nu}_{a}^{2}\wp(z+\omega_{a})-\nu_{a}^{2}\wp(u+\omega_{a})\Big{)}\,,}\end{array}$ (3.17)
where similarly to (A.21)
$\begin{array}[]{c}\left(\begin{array}[]{c}\bar{\nu}_{0}\\ \bar{\nu}_{1}\\ \bar{\nu}_{2}\\ \bar{\nu}_{3}\end{array}\right)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{cccc}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right)\left(\begin{array}[]{c}\nu_{0}\\ \nu_{1}\\ \nu_{2}\\ \nu_{3}\end{array}\right)\,.\end{array}$ (3.18)
From (A.21) we conclude that
$\begin{array}[]{c}\displaystyle{v(z,u|\nu)=v(u,z|\bar{\nu})=\sum\limits_{a=0}^{3}\bar{\nu}_{a}\varphi_{a}(2u,z+\omega_{a})\,.}\end{array}$ (3.19)
By differentiating (3.17) with respect to $u$ we get:
$\begin{array}[]{c}\displaystyle{v(z,u)v^{\prime}(z,-u)-v(z,-u)v^{\prime}(z,u)=\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(u+\omega_{a})\,.}\end{array}$ (3.20)
The calculation of the Hamiltonian (1.11) is standard. Using (A.16) and (3.17) one gets
$\begin{array}[]{c}\displaystyle{\frac{1}{4}\,{\rm tr}(L^{2}(z))=H+n(n-1)g^{2}\wp(z)+\frac{n}{2}\sum\limits_{a=0}^{3}\bar{\nu}_{a}^{2}\wp(z+\omega_{a})\,.}\end{array}$ (3.21)
In order to prove the Lax equation for the Lax pair (3.2) it is helpful to use the following identities:
$\begin{array}[]{c}\displaystyle{v(x,u)\phi(x+y,w-u)+v(x,w)\phi(x-y,u-w)+v(y,-u)\phi(x+y,u+w)=}\\ \\ \displaystyle{=v(y,w)\phi(x-y,u+w)}\end{array}$ (3.22)
and
$\begin{array}[]{c}\phi(z,u-w)(v^{\prime}(0,w)-v^{\prime}(0,u))=\\ \\ 2v(-z,w)f(z,u+w)+2v(z,u)f(-z,u+w)+v^{\prime}(-z,w)\phi(z,u+w)+v^{\prime}(z,u)\phi(-z,u+w)\,,\end{array}$ (3.23)
where
$\begin{array}[]{c}\displaystyle{v^{\prime}(0,u)=\sum\limits_{a=0}^{3}\nu_{a}f(0,u+\omega_{a})=-\sum\limits_{a=0}^{3}\nu_{a}E_{2}(u+\omega_{a})=-\sum\limits_{a=0}^{3}\nu_{a}\wp(u+\omega_{a})+\frac{\vartheta^{\prime\prime\prime}(0)}{3\vartheta^{\prime}(0)}\sum\limits_{a=0}^{3}\nu_{a}\,.}\end{array}$ (3.24)
The latter relation is derived from (A.13) and (A.15), and the identities (3.22) and (3.23) are proved in the Appendix.
3.2 Lax equation
Let us verify the Lax equation (1.4) in detail in order to generalize it in the Section 5. We are going to use the identity (3.23) in the form:
$\begin{array}[]{c}\displaystyle{\phi(z,q_{ij}^{+})v^{\prime}(z,-q_{j})-v^{\prime}(z,q_{i})\phi(z,-q_{ij}^{+})+2v(z,q_{i})f(z,-q_{ij}^{+})-2f(z,q_{ij}^{+})v(z,-q_{j})=}\\ \\ \displaystyle{=\phi(z,q_{ij})(v^{\prime}(0,q_{j})-v^{\prime}(0,q_{i}))\,.}\end{array}$ (3.25)
It is natural to subdivide verification by considering four $n\times n$ matrix blocks separately.
The block 11.
We need to show that
$\dot{L}^{11}(z)=L^{11}(z)M^{11}(z)-M^{11}(z)L^{11}(z)+L^{12}(z)M^{21}(z)-M^{12}(z)L^{21}(z)\,.$ (3.26)
For the diagonal part we have:
$\begin{array}[]{c}\displaystyle{\dot{L^{11}_{ii}}=L^{12}_{ii}M^{21}_{ii}-M^{12}_{ii}L^{21}_{ii}+\sum\limits_{k:k\neq i}^{n}\Big{(}L^{12}_{ik}M^{21}_{ki}-M^{12}_{ik}L^{21}_{ki}+L^{11}_{ik}M^{11}_{ki}-M^{11}_{ik}L^{11}_{ki}\Big{)}\,,}\end{array}$ (3.27)
or, explicitly:
$\begin{array}[]{c}\displaystyle{\dot{p_{i}}=\frac{1}{2}v(z,q_{i})v^{\prime}(z,-q_{i})-\frac{1}{2}v(z,-q_{i})v^{\prime}(z,q_{i})+}\\ \\ \displaystyle{+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}\phi(z,q^{+}_{ik})f(z,-q^{+}_{ik})-\phi(z,-q^{+}_{ik})f(z,q^{+}_{ik})+}\\ \\ \displaystyle{+\phi(z,q_{ik})f(z,q_{ki})-\phi(z,q_{ki})f(z,q_{ik})\Big{)}=}\\ \\ \displaystyle{=\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(q_{i}+\omega_{a})+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}\wp^{\prime}(q^{+}_{ik})+\wp^{\prime}(q_{ik})\Big{)}\,,}\end{array}$ (3.28)
where we used (3.20) and (2.3).
For the non-diagonal part we have:
$\begin{array}[]{c}\displaystyle{\dot{L^{11}_{ij}}=L^{12}_{ij}M^{21}_{jj}-M^{12}_{ij}L^{21}_{jj}+L^{11}_{ij}M^{11}_{jj}-M^{11}_{ij}L^{11}_{jj}+L^{12}_{ii}M^{21}_{ij}-M^{12}_{ii}L^{21}_{ij}+L^{11}_{ii}M^{11}_{ij}-M^{11}_{ii}L^{11}_{ij}+}\\ \\ \displaystyle{+\sum\limits_{k:k\neq i,j}^{n}\Big{(}L^{12}_{ik}M^{21}_{kj}-M^{12}_{ik}L^{21}_{kj}+L^{11}_{ik}M^{11}_{kj}-M^{11}_{ik}L^{11}_{kj}\Big{)}\,.}\end{array}$ (3.29)
The l.h.s. equals $(\dot{q_{i}}-\dot{q_{j}})f(z,q_{ij})$. The r.h.s is
$\begin{array}[]{c}\displaystyle{f(z,q_{ij})(p_{i}-p_{j})+\underline{\underline{\frac{1}{2}\,\phi(z,q^{+}_{ij})v^{\prime}(z,-q_{j})+f(z,q^{+}_{ij})v(-z,q_{j})+}}}\\ \\ \displaystyle{\underline{\underline{+\frac{1}{2}\,\phi(-z,q^{+}_{ij})v^{\prime}(z,q_{i})+f(z,-q^{+}_{ij})v(z,q_{i})+}}}\\ \\ \displaystyle{+\underline{g\sum\limits_{k:k\neq i,j}^{n}\phi(z,q_{ij})\Big{(}\wp(q_{jk})+\wp(q^{+}_{jk})-\wp(q_{ik})-\wp(q^{+}_{ik})\Big{)}}+}\\ \\ \displaystyle{\underline{\underline{+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}\phi(z,q_{ij})\Big{(}\wp(q_{j}+\omega_{a})-\wp(q_{i}+\omega_{a})\Big{)}+}}}\\ \\ \displaystyle{+\underline{g\sum\limits_{k:k\neq i,j}^{n}\Big{(}\phi(z,q^{+}_{ik})f(z,-q^{+}_{kj})-\phi(z,-q^{+}_{kj})f(z,q^{+}_{ki})+\phi(z,q_{ik})f(z,q_{kj})-\phi(z,q_{kj})f(z,q_{ik})\Big{)}}=}\\ \\ \displaystyle{=f(z,q_{ij})(p_{i}-p_{j})\,,}\end{array}$ (3.30)
where the underlined and the double underlined terms are cancelled out respectively due to (A.15) and (3.23), (3.24).
The block 12.
Here
$\dot{L}^{12}(z)=L^{11}(z)M^{12}(z)+L^{12}(z)M^{22}(z)-M^{11}(z)L^{12}(z)-M^{12}(z)L^{22}(z)\,.$ (3.31)
For the diagonal part we have:
$\begin{array}[]{c}\displaystyle{\dot{L^{12}_{ii}}=L^{11}_{ii}M^{12}_{ii}-M^{12}_{ii}L^{22}_{ii}+\sum\limits_{k:k\neq i}^{n}\Big{(}L^{11}_{ik}M^{12}_{ki}-M^{11}_{ik}L^{12}_{ki}+L^{12}_{ik}M^{22}_{ki}-M^{12}_{ik}L^{22}_{ki}\Big{)}\,,}\end{array}$ (3.32)
that is
$\begin{array}[]{c}\displaystyle{\dot{q_{i}}v^{\prime}(z,q_{i})=p_{i}v^{\prime}(z,q_{i})+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}\phi(z,q_{ik})f(z,q^{+}_{ki})-f(z,q_{ik})\phi(z,q^{+}_{ki})+}\\ \\ \displaystyle{\phi(z,q^{+}_{ik})f(z,q_{ik})-f(z,q^{+}_{ik})\phi(z,q_{ik})\Big{)}=p_{i}v^{\prime}(z,q_{i})\,.}\end{array}$ (3.33)
For the non-diagonal part we have:
$\begin{array}[]{c}\displaystyle{\dot{L^{12}_{ij}}=L^{11}_{ij}M^{12}_{jj}-M^{11}_{ij}L^{12}_{jj}+L^{11}_{ii}M^{12}_{ij}-M^{11}_{ii}L^{12}_{ij}+L^{12}_{ii}M^{22}_{ij}-M^{12}_{ii}L^{22}_{ij}+L^{12}_{ij}M^{22}_{jj}-M^{12}_{ij}L^{22}_{jj}+}\\ \\ \displaystyle{+\sum\limits_{k:k\neq i,j}^{n}\Big{(}L^{11}_{ik}M^{12}_{kj}-M^{11}_{ik}L^{12}_{kj}+L^{12}_{ik}M^{22}_{kj}-M^{12}_{ik}L^{22}_{kj}\Big{)}\,.}\end{array}$ (3.34)
The explicit form can be obtained from (3.30) by replacing $q_{j},p_{j}$ to $-q_{j},-p_{j}$.
The block 21 and 22
For other blocks we need to show that
$\dot{L}^{21}(z)=L^{21}(z)M^{11}(z)+L^{22}(z)M^{21}(z)-M^{21}(z)L^{11}(z)-M^{22}(z)L^{21}(z),$ (3.35)
$\dot{L}^{22}(z)=L^{21}(z)M^{12}(z)+L^{22}(z)M^{22}(z)-M^{21}(z)L^{12}(z)-M^{22}(z)L^{22}(z),$ (3.36)
which follow from (3.31) and (3.26) respectively due to the properties (3.12) and (3.13).
4 Shibukawa–Ueno elliptic $R$-operator and $K$-matrices
4.1 Associative Yang-Baxter equation and $K$-operators
In [29] Y. Shibukawa and K.Ueno introduced an elliptic operator $R_{ij}$:
$R^{\hbar}_{ij}(q)=\phi(\hbar,x_{i}-x_{j})-\phi(q,x_{i}-x_{j})\hat{s}_{ij}\,,$ (4.1)
which satisfies the quantum Yang-Baxter equation
$R^{\hbar}_{ij}(q_{i}-q_{j})R^{\hbar}_{ik}(q_{i}-q_{k})R^{\hbar}_{jk}(q_{j}-q_{k})=R^{\hbar}_{jk}(q_{j}-q_{k})R^{\hbar}_{ik}(q_{i}-q_{k})R^{\hbar}_{ij}(q_{i}-q_{j}).$ (4.2)
Here the symmetric group $S_{N}$ acts by exchanging the variables $x_{i}$ and $\hat{s}_{ij}$ denotes the corresponding transposition
$\hat{s}_{ij}f(\dots,x_{i},\dots,x_{j},\dots)=f(\dots,x_{j},\dots,x_{i},\dots).$ (4.3)
In [17] Y. Komori and K. Hikami proposed the following boundary $K$-operator:
$K^{\hbar}_{i}(q)\equiv K^{\hbar}_{i}(q|\bar{\nu})=v(q,x_{i}|\bar{\nu})-v(\hbar,x_{i}|\bar{\nu})\hat{t}_{i}\,,$ (4.4)
where $\hat{t}_{i}$ is a reflection operator
$\hat{t}_{i}f(x_{1}\dots,x_{i},\dots,x_{n})=f(x_{1}\dots,-x_{i},\dots,x_{n})\,.$ (4.5)
The operator relations are as follows:
$\hat{t}_{i}^{2}=1,\qquad\hat{s}_{ij}\hat{t}_{j}=\hat{t}_{i}\hat{s}_{ij}\,.$ (4.6)
It was also shown in [17] that the reflection equation holds true for these operators, that is
$R_{ij}^{\hbar}(q_{i}-q_{j})K_{i}^{\hbar}(q_{i})R_{ji}^{\hbar}(q_{i}+q_{j})K_{j}^{\hbar}(q_{j})=K_{j}^{\hbar}(q_{j})R_{ij}^{\hbar}(q_{i}+q_{j})K_{i}^{\hbar}(q_{i})R_{ji}^{\hbar}(q_{i}-q_{j})\,.$ (4.7)
Let us introduce the following operator:
$\tilde{R}_{ij}^{\hbar}(q)=\phi(\hbar,x_{i}+x_{j})-\phi(q,x_{i}+x_{j})\hat{s}_{ij}\hat{t}_{i}\hat{t}_{j}\,,$ (4.8)
where the operators $\tilde{R}_{ij}^{\hbar}(q)$ and $R_{ij}^{\hbar}(q)$ are connected by conjugation by $\hat{t}_{j}$:
$\tilde{R}_{ij}^{\hbar}(q)=\hat{t}_{j}R_{ij}^{\hbar}(q)\hat{t}_{j}\,.$ (4.9)
Conjugating (4.2) by $\hat{t}_{k}$ ad using (4.6) we have the following equation:
$R^{\hbar}_{ij}(q_{i}-q_{j})\tilde{R}^{\hbar}_{ik}(q_{i}-q_{k})\tilde{R}^{\hbar}_{jk}(q_{j}-q_{k})=\tilde{R}^{\hbar}_{jk}(q_{j}-q_{k})\tilde{R}^{\hbar}_{ik}(q_{i}-q_{k})R^{\hbar}_{ij}(q_{i}-q_{j}).$ (4.10)
Also, we have:
$\tilde{R}_{ij}^{\hbar}(q)=-\hat{t}_{i}R_{ij}^{-\hbar}(-q)\hat{t}_{i}\,.$ (4.11)
Introduce another one operator $\tilde{K}$:
$\tilde{K}^{\hbar}_{i}(q)\equiv\tilde{K}^{\hbar}_{i}(q|\bar{\nu})=v(\hbar,x_{i}|\bar{\nu})-v(q,x_{i}|\bar{\nu})\hat{t}_{i}\,.$ (4.12)
It is related to the operator (4.4) as
$\tilde{K}^{\hbar}_{i}(q)=-{K}^{\hbar}_{i}(q)\hat{t}_{i}=\hat{t}_{i}K^{-\hbar}_{i}(-q)\,.$ (4.13)
Using (4.9), (4.11) and (4.13), we can rewrite the reflection equation (4.7) in terms of $\tilde{K}^{\hbar}_{i}(q)$:
$R_{ij}^{\hbar}(q_{i}-q_{j})\tilde{K}_{i}^{\hbar}(q_{i})\tilde{R}_{ij}^{\hbar}(q_{i}+q_{j})\tilde{K}_{j}^{\hbar}(q_{j})=\tilde{K}_{j}^{\hbar}(q_{j})\tilde{R}_{ij}^{\hbar}(q_{i}+q_{j})\tilde{K}_{i}^{\hbar}(q_{i})R_{ij}^{\hbar}(q_{i}-q_{j})\,.$ (4.14)
Lemma 4.1
The following unitarity properties hold true:
$R_{ij}^{\hbar}(q)R_{ji}^{\hbar}(-q)=(\wp(\hbar)-\wp(q))\,,$ (4.15)
$\tilde{R}_{ij}^{\hbar}(q)\tilde{R}_{ji}^{-\hbar}(q)=(\wp(q)-\wp(\hbar))\,,$ (4.16)
$K^{\hbar}_{i}(q)K^{\hbar}_{i}(-q)=\sum_{a=0}^{3}\nu^{2}_{a}\left(\wp(\hbar+\omega_{a})-\wp(q+\omega_{a})\right)\,,$ (4.17)
$\tilde{K}^{\hbar}_{i}(q)\tilde{K}^{-\hbar}_{i}(q)=\sum_{a=0}^{3}\nu^{2}_{a}\left(\wp(q+\omega_{a})-\wp(\hbar+\omega_{a})\right)\,.$ (4.18)
Proof. Let us prove (4.15):
$\begin{array}[]{ccc}R_{ij}^{\hbar}(q)R_{ji}^{\hbar}(-q)=\left(\phi(\hbar,x_{i}-x_{j})-\phi(q,x_{i}-x_{j})s_{ij}\right)\left(\phi(\hbar,x_{j}-x_{i})-\phi(-q,x_{j}-x_{i})s_{ij}\right)=\\ \\ =\phi(\hbar,x_{i}-x_{j})\phi(\hbar,x_{j}-x_{i})-\phi(q,x_{i}-x_{j})\phi(\hbar,x_{i}-x_{j})s_{ij}\\ \\ -\phi(\hbar,x_{i}-x_{j})\phi(-q,x_{j}-x_{i})s_{ij}+\phi(q,x_{i}-x_{j})\phi(-q,x_{i}-x_{j})=\\ \\ =\phi(\hbar,x_{i}-x_{j})\phi(\hbar,x_{j}-x_{i})-\phi(q,x_{i}-x_{j})\phi(q,x_{j}-x_{i})=\wp(\hbar)-\wp(q).\end{array}$ (4.19)
In the first equality we used $s_{ij}^{2}=1$, then (A.37), and (A.16) in the last equality.
The relation (4.16) follows form (4.15) and (4.9), (4.11), (4.6). Indeed,
$\hat{t}_{i}R_{ij}^{\hbar}(q)R_{ji}^{\hbar}(-q)\hat{t_{i}}=\hat{t}_{i}R_{ij}^{\hbar}(q)(\hat{t}_{i})^{2}R_{ji}^{\hbar}(-q)\hat{t_{i}}=-\tilde{R}_{ij}^{\hbar}(q)\tilde{R}_{ji}^{-\hbar}(q).$ (4.20)
Next, let us prove (4.17):
$\begin{array}[]{ccc}K^{\hbar}_{i}(q)K^{\hbar}_{i}(-q)=\left(v(q,x_{i})-v(\hbar,x_{i})\hat{t}_{i}\right)\left(v(-q,x_{i})-v(\hbar,x_{i})\hat{t}_{i}\right)=\\ \\ =v(q,x_{i})v(-q,x_{i})-v(q,x_{i})v(\hbar,x_{i})\hat{t}_{i}-v(\hbar,x_{i})v(-q,-x_{i})\hat{t}_{i}+v(\hbar,x_{i})v(\hbar,-x_{i})=\\ \\ \displaystyle=v(q,x_{i})v(-q,x_{i})+v(\hbar,x_{i})v(\hbar,-x_{i})=\end{array}$ (4.21)
$\begin{array}[]{ccc}\displaystyle=\sum_{a=0}^{3}\bar{\nu}^{2}_{a}\wp(x_{i}+\omega_{a})-\sum_{a=0}^{3}\nu^{2}_{a}\wp(q+\omega_{a})+\sum_{a=0}^{3}\nu^{2}_{a}\wp(\hbar+\omega_{a})-\sum_{a=0}^{3}\bar{\nu}^{2}_{a}\wp(x_{i}+\omega_{a})=\\ \displaystyle=\sum_{a=0}^{3}\nu^{2}_{a}\left(\wp(\hbar+\omega_{a})-\wp(q+\omega_{a})\right).\end{array}$
The antisymmetry property (A.37) was used in the second equality, and in the third one equality we used (A.31) and (3.17).
In order to show (4.18) one should use the relations (4.6) and (4.13):
$K^{\hbar}_{i}(q)K^{\hbar}_{i}(-q)=K^{\hbar}_{i}(q)(\hat{t}_{i})^{2}K^{\hbar}_{i}(-q)=-\tilde{K}^{\hbar}_{i}(q)\tilde{K}_{i}^{-\hbar}(q)\,.$ (4.22)
[…]
Let us introduce notations $F_{ij}^{z}(q)$ (c.f. (2.8)) and $\tilde{F}_{ij}^{z}(q)$ for the corresponding partial derivatives of $R_{ij}^{z}(q)$ and $\tilde{R}_{ij}^{z}(q)$:
$F_{ij}^{z}(q):=\partial_{q}R_{ij}^{z}(q)\qquad\tilde{F}_{ij}^{z}(q):=\partial_{q}\tilde{R}_{ij}^{z}(q)\,,$ (4.23)
and denote by $Y^{\hbar}_{i}(q)$ and $\tilde{Y}^{\hbar}_{i}(q)$ the corresponding partial derivatives of the reflection operators:
$Y^{\hbar}_{i}(q):=\partial_{q}K^{\hbar}_{i}(q|\bar{\nu})\qquad\tilde{Y}^{\hbar}_{i}(q):=\partial_{q}\tilde{K}^{\hbar}_{i}(q|\bar{\nu})\,.$ (4.24)
Corollary 4.1
The following identities hold true :
$R_{ij}^{\hbar}(q)F_{ji}^{\hbar}(-q)-F_{ij}^{\hbar}(q)R_{ji}^{\hbar}(-q)=\wp^{\prime}(q)\,,$ (4.25)
$\tilde{R}_{ij}^{\hbar}(q)\tilde{F}_{ji}^{-\hbar}(q)+\tilde{F}_{ij}^{\hbar}(q)\tilde{R}_{ji}^{-\hbar}(q)=\wp^{\prime}(q)\,,$ (4.26)
$K^{\hbar}_{i}(q)Y^{\hbar}_{i}(-q)-Y^{\hbar}_{i}(q)K^{\hbar}_{i}(-q)=\sum_{a=0}^{3}\nu^{2}_{a}\wp^{\prime}(q+\omega_{a})\,,$ (4.27)
$\tilde{Y}^{\hbar}_{i}(q)\tilde{K}^{-\hbar}_{i}(q)+\tilde{K}^{\hbar}_{i}(q)\tilde{Y}^{-\hbar}_{i}(q)=\sum_{a=0}^{3}\nu^{2}_{a}\wp^{\prime}(q+\omega_{a})\,.$ (4.28)
[…]
4.2 ${\rm BC}_{n}$ associative Yang-Baxter equations with parameters
The construction presented below is motivated by the Kirillov’s ${\rm B}$-type associative Yang-Baxter equation, see Section 5 in [18]. The associative Yang-Baxter equation corresponding to ${\rm A}$-type root systems (1.8) in the absence of parameters takes the following simple form:
$\begin{array}[]{c}r_{ij}r_{jk}=r_{ik}r_{ij}+r_{jk}r_{ik}\qquad\text{ for distinct }i,j,k.\end{array}$ (4.29)
In order to define the associative classical Yang-Baxter algebra of type ${\rm A}_{n}$ generated by the elements $r_{ij}$ for $1\leq i\neq j\leq n$, we extend (4.29) with the antisymmetry condition and commutativity relation:
$\begin{array}[]{c}{\displaystyle r_{ij}r_{kl}=r_{kl}r_{ij}\qquad\text{ for distinct }i,j,k,l,}\\ \\ {\displaystyle r_{ij}=-r_{ji}}\,.\end{array}$ (4.30)
The equation (1.8) with parameters was suggested by A. Polishchuk in [27]. The relation (4.29) looks as the half of the classical Yang-Baxter equation. From (4.29) and (4.30) the classical Yang-Baxter equation follows:
$[r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0.$ (4.31)
The classical associative Yang-Baxter algebra of type ${\rm B}_{n}$ was proposed in [18]. It is generated by the elements $r_{ij}$, $\tilde{r}_{ij}$ and $y_{i}$ for $1\leq i\neq j\leq n$ satisfying the set of defining relations:
  • symmetry and antisymmetry relations for $i\neq j$:
    $r_{ij}=-r_{ji}\,,\qquad\tilde{r}_{ij}=\tilde{r}_{ji}\,,$ (4.32)
  • permutation relations for distinct $i,j,k,l$:
    $\begin{array}[]{c}{\displaystyle y_{i}y_{j}=y_{j}y_{i}\,,\qquad y_{i}r_{kl}=r_{kl}y_{i}\,,\qquad y_{i}r_{kl}=r_{kl}y_{i}\,,}\\ \\ {\displaystyle r_{ij}r_{kl}=r_{kl}r_{ij},\qquad\tilde{r}_{ij}r_{kl}=r_{kl}\tilde{r}_{ij},\qquad\tilde{r}_{ij}\tilde{r}_{kl}=\tilde{r}_{kl}\tilde{r}_{ij}\,,}\end{array}$ (4.33)
  • three term relations for distinct $i,j,k$:
    $\begin{array}[]{c}{\displaystyle r_{ij}r_{jk}=r_{ik}r_{ij}+r_{jk}r_{ik}\,,\qquad r_{ij}\tilde{r}_{jk}=r_{ik}\tilde{r}_{ij}+\tilde{r}_{jk}r_{ik}\,,}\\ \\ {\displaystyle r_{jk}\tilde{r}_{ik}=\tilde{r}_{ij}r_{jk}+\tilde{r}_{ik}\tilde{r}_{ij}\,,\qquad\tilde{r}_{jk}r_{ik}=\tilde{r}_{ij}\tilde{r}_{jk}+\tilde{r}_{ik}\tilde{r}_{ij},}\end{array}$ (4.34)
  • four term relation for $i\neq j$:
    $r_{ij}y_{j}=y_{i}r_{ij}+\tilde{r}_{ij}y_{i}+y_{j}\tilde{r}_{ij}\,.$ (4.35)
Proposition 4.1
The representation of the classical associative Yang-Baxter algebra of type ${\rm B}_{n}$ can be realized by the following operators:
$r_{ij}:=\phi(q_{i}-q_{j},x_{i}-x_{j})\hat{s}_{ij},\qquad\tilde{r}_{ij}:=\phi(q_{i}+q_{j},x_{i}+x_{j})\hat{s}_{ij}\hat{t}_{i}\hat{t}_{j},\qquad y_{i}:=v(q_{i},x_{i})\hat{t}_{i}\,.$ (4.36)
[…]
Below we formulate a deformation of the relations (4.32), (4.33), (4.34) and (4.35) in the presence of parameters.
Proposition 4.2
The ${\rm BC}_{n}$ associative Yang-Baxter equation with parameters: the following set of relations for the operators (4.1), (4.8), (4.12) hold true:
  • Symmetry and antisymmetry relations:
    $\tilde{R}^{z}_{ij}(q)=\tilde{R}^{z}_{ji}(q)\,,$ (4.37)
    $R^{z}_{ij}(q)=-R_{ji}^{-z}(-q)\,,$ (4.38)
  • Permutation relation for distinct $i,j,k,l$:
    $\tilde{K}^{z}_{i}(q_{i})\tilde{K}^{w}_{j}(q_{j})=\tilde{K}^{w}_{j}(q_{j})\tilde{K}_{i}^{z}(q_{i})\,,\qquad\tilde{K}^{z}_{i}(q_{i})R^{w}_{kl}(q_{kl})=R^{w}_{kl}(q_{kl})\tilde{K}_{i}^{z}(q_{i})\,,$ (4.39)
    $\tilde{K}^{z}_{i}(q_{i})\tilde{R}^{w}_{kl}(q_{kl})=\tilde{R}^{w}_{kl}(q_{kl})\tilde{K}_{i}^{z}(q_{i})\,,\qquad R^{z}_{ij}(q_{ij})R^{w}_{kl}(q_{kl})=R^{w}_{kl}(q_{kl})R^{z}_{ij}(q_{ij})\,,$ (4.40)
    $R^{z}_{ij}(q_{ij})\tilde{R}^{w}_{kl}(q_{kl})=\tilde{R}^{w}_{kl}(q_{kl})R^{z}_{ij}(q_{ij})\,,\qquad\tilde{R}^{z}_{ij}(q_{ij})\tilde{R}^{w}_{kl}(q_{kl})=\tilde{R}^{w}_{kl}(q_{kl})\tilde{R}^{z}_{ij}(q_{ij})\,,$ (4.41)
  • Three term relations:
    $R_{ij}^{z}(q_{i}-q_{j})R_{jk}^{w}(q_{j}-q_{k})=R_{ik}^{w}(q_{i}-q_{k})R_{ij}^{z-w}(q_{i}-q_{j})+R_{jk}^{w-z}(q_{j}-q_{k})R_{ik}^{z}(q_{i}-q_{k})\,,$ (4.42)
    $R_{ij}^{z}(q_{i}-q_{j})\tilde{R}_{jk}^{w}(q_{j}+q_{k})=\tilde{R}_{ik}^{w}(q_{i}+q_{k})R_{ij}^{z-w}(q_{i}-q_{j})+\tilde{R}_{jk}^{w-z}(q_{j}+q_{k})\tilde{R}_{ik}^{z}(q_{i}+q_{k})\,,$ (4.43)
    $R_{jk}^{z}(q_{j}-q_{k})\tilde{R}_{ik}^{w}(q_{i}+q_{k})=\tilde{R}_{ij}^{w}(q_{i}+q_{j})R_{jk}^{z-w}(q_{j}-q_{k})+\tilde{R}_{ik}^{w-z}(q_{i}+q_{k})\tilde{R}_{ij}^{z}(q_{i}+q_{j})\,,$ (4.44)
    $\tilde{R}_{jk}^{z}(q_{j}+q_{k})R_{ik}^{w}(q_{i}-q_{k})=\tilde{R}_{ij}^{w}(q_{i}+q_{j})\tilde{R}_{jk}^{z-w}(q_{j}+q_{k})+R_{ik}^{w-z}(q_{i}-q_{k})\tilde{R}_{ij}^{z}(q_{i}+q_{j})\,,$ (4.45)
  • The four term relation (generalization of (3.22)):
    $R_{ij}^{w+z}(q_{i}-q_{j})\tilde{K}^{w}_{j}(q_{j})=\tilde{K}^{w}_{i}(q_{i})R_{ij}^{z-w}(q_{i}-q_{j})+\tilde{R}_{ij}^{w-z}(q_{i}+q_{j})\tilde{K}^{z}_{i}(q_{i})+\tilde{K}_{j}^{-z}(q_{j})\tilde{R}_{ij}^{w+z}(q_{i}+q_{j})\,.$ (4.46)
Proof. The relations (4.37) and (4.38) follows from (A.37). The relations (4.39), (4.40) and (4.41) are valid since $t_{i}$ , $\hat{t}_{j}$ and $s_{kl}$ commute for distinct $i,j,k,l$ and do not change any arguments in the corresponding functions $v(z,x_{i})$, $v(w,x_{j})$ and $\phi(w,x_{k}\pm x_{l})$. Also $s_{ij}s_{kl}=s_{kl}s_{ij}$ and $s_{kl}\phi(z,x_{i}\pm x_{j})=\phi(z,x_{i}\pm x_{j})s_{kl}$.
Let us prove (4.42). We use the following short notations $q_{ij}=q_{i}-q_{j}$ and $x_{ij}=x_{i}-x_{j}$ and calculate each summand in (4.42):
$\begin{array}[]{ccc}R_{ij}^{z}(q_{i}-q_{j})R_{jk}^{w}(q_{j}-q_{k})=\left(\phi(z,x_{ij})-\phi(q_{ij},x_{ij})s_{ij}\right)\left(\phi(w,x_{jk})-\phi(q_{jk},x_{jk})s_{jk}\right)=\\ \\ =\phi(z,x_{ij})\phi(w,x_{jk})\uuline{-\phi(z,x_{ij})\phi(q_{jk},x_{jk})s_{jk}}\\ \\ \uwave{-\phi(q_{ij},x_{ij})\phi(w,x_{ik})s_{ij}}+\phi(q_{ij},x_{ij})\phi(q_{jk},x_{ik})s_{ij}s_{jk}\,,\end{array}$ (4.47)
$\begin{array}[]{ccc}R_{ik}^{w}(q_{i}-q_{k})R_{ij}^{z-w}(q_{i}-q_{j})=\left(\phi(w,x_{ik})-\phi(q_{ik},x_{ik})s_{ik}\right)\left(\phi(z-w,x_{ij})-\phi(q_{ij},x_{ij})s_{ij}\right)=\\ \\ =\phi(w,x_{ik})\phi(z-w,x_{ij})\uwave{-\phi(w,x_{ik})\phi(q_{ij},x_{ij})s_{ij}}\\ \\ \underline{-\phi(q_{ik},x_{ik})\phi(z-w,x_{kj})s_{ik}}+\phi(q_{ik},x_{ik})\phi(q_{ij},x_{kj})s_{ik}s_{ij}\,,\end{array}$ (4.48)
$\begin{array}[]{ccc}R_{jk}^{w-z}(q_{j}-q_{k})R_{ik}^{z}(q_{i}-q_{k})=\left(\phi(w-z,x_{jk})-\phi(q_{jk},x_{jk})s_{jk}\right)\left(\phi(z,x_{ik})-\phi(q_{ik},x_{ik})s_{ik}\right)=\\ \\ =\phi(w-z,x_{jk})\phi(z,x_{ik})\underline{-\phi(w-z,x_{jk})\phi(q_{ik},x_{ik})s_{ik}}\\ \\ \uuline{-\phi(q_{jk},x_{jk})\phi(z,x_{ij})s_{jk}}+\phi(q_{jk},x_{jk})\phi(q_{ik},x_{ij})s_{jk}s_{ik}\,.\end{array}$ (4.49)
The corresponding underlined terms are cancelled identically due to the antisymmetry property (A.37). The rest of the terms vanish due to the Fay identity (A.12) written as
$\phi(z,x_{ij})\phi(w,x_{jk})-\phi(w,x_{ik})\phi(z-w,x_{ij})-\phi(w-z,x_{jk})\phi(z,x_{ik})=0$ (4.50)
and
$\begin{array}[]{ccc}\phi(q_{ij},x_{ij})\phi(q_{jk},x_{ik})-\phi(q_{ik},x_{ik})\phi(q_{ij},x_{kj})-\phi(q_{jk},x_{jk})\phi(q_{ik},x_{ij})=\\ \\ \,=\phi(q_{ij},x_{ij})\phi(q_{jk},x_{ik})+\phi(q_{ik},x_{ik})\phi(q_{ji},x_{jk})-\phi(q_{jk},x_{jk})\phi(q_{ik},x_{ij})=0\end{array}$ (4.51)
and the following equality in the symmetric group $S_{N}$: $s_{ij}s_{jk}=s_{ik}s_{ij}=s_{jk}s_{ik}$.
To prove (4.43) we conjugate (4.42) by the operator $\hat{t}_{j}$ and use that $\hat{t}_{j}^{2}=1$ and $\hat{t}_{j}$ commutes with $R_{ik}^{\hbar}(q_{i}-q_{k})$:
$\hat{t}_{j}R_{ij}^{z}(q_{i}-q_{j})\hat{t}_{j}^{2}R_{jk}^{w}(q_{j}-q_{k})\hat{t}_{j}=R_{ik}^{w}(q_{i}-q_{k})\hat{t}_{j}R_{ij}^{z-w}(q_{i}-q_{j})\hat{t}_{j}+\hat{t}_{j}R_{jk}^{w-z}(q_{j}-q_{k})\hat{t}_{j}R_{ik}^{z}(q_{i}-q_{k})\,,$ (4.52)
then using (4.9) we rewrite (4.52) as
$\tilde{R}_{ij}^{z}(q_{i}-q_{j})\tilde{R}_{jk}^{w}(q_{j}-q_{k})=R_{ik}^{w}(q_{i}-q_{k})\tilde{R}_{ij}^{z-w}(q_{i}-q_{j})+\tilde{R}_{jk}^{w-z}(q_{j}-q_{k})R_{ik}^{z}(q_{i}-q_{k})\,.$ (4.53)
By replacing $q_{k}\rightarrow-q_{k}$ we obtain (4.43). In the same manner one can prove (4.44) and (4.45) conjugating (4.42) by $\hat{t}_{k}$ and $\hat{t}_{i}$, respectively, and using (4.11).
To prove (4.46) we calculate each summand moving all operators $\hat{s}_{ij}$, $t_{i}$ and $t_{j}$ to the right and using (4.6):
$R_{ij}^{w+z}(q_{i}-q_{j})\tilde{K}^{w}_{j}(q_{j})=$ $\phi(w+z,x_{i}-x_{j})v(w,x_{j})\underline{-\phi(q_{i}-q_{j},x_{i}-x_{j})v(w,x_{i})\hat{s}_{ij}}$
$\uuline{-\phi(z+w,x_{i}-x_{j})v(q_{j},x_{j})\hat{t}_{j}}+\phi(q_{i}-q_{j},x_{i}-x_{j})v(q_{j},x_{i})\hat{s}_{ij}\hat{t}_{j}\,,$
$\tilde{K}^{w}_{i}(q_{i})R_{ij}^{z-w}(q_{i}-q_{j})=$ $v(w,x_{i})\phi(z-w,x_{i}-x_{j})\underline{-v(w,x_{i})\phi(q_{i}-q_{j},x_{i}-x_{j})\hat{s}_{ij}}$
$\uwave{-v(q_{i},x_{i})\phi(z-w,-x_{i}-x_{j})\hat{t}_{i}}+v(q_{i},x_{i})\phi(q_{i}-q_{j},-x_{i}-x_{j})\hat{s}_{ij}\hat{t}_{j}\,,$
$\tilde{R}_{ij}^{w-z}(q_{i}+q_{j})\tilde{K}^{z}_{i}(q_{i})=$ $\phi(w-z,x_{i}+x_{j})v(z,x_{i})\dashuline{-\phi(q_{i}+q_{j},x_{i}+x_{j})v(z,-x_{j})\hat{s}_{ij}\hat{t}_{i}\hat{t}_{j}}$
$\uwave{-\phi(w-z,x_{i}+x_{j})v(q_{i},x_{i})\hat{t}_{i}}+\phi(q_{i}+q_{j},x_{i}+x_{j})v(q_{i},-x_{j})\hat{s}_{ij}\hat{t}_{j}\,,$
$\tilde{K}_{j}^{-z}(q_{j})\tilde{R}_{ij}^{w+z}(q_{i}+q_{j})=$ $v(-z,x_{j})\phi(w+z,x_{i}+x_{j})\dashuline{-v(-z,x_{j})\phi(q_{i}+q_{j},x_{i}+x_{j})\hat{s}_{ij}\hat{t}_{i}\hat{t}_{j}}$
$\uuline{-v(q_{j},x_{j})\phi(w+z,x_{i}-x_{j})\hat{t}_{j}}+v(q_{j},x_{j})\phi(q_{i}+q_{j},x_{i}-x_{j})\hat{s}_{ij}\hat{t}_{j}\,.$
The corresponding underlined terms are cancelled identically due (A.37). To finish the proof one should check the identities:
$v(w,x_{j})\phi(w+z,x_{i}-x_{j})=v(w,x_{i})\phi(z-w,x_{i}-x_{j})$ $+$ (4.58)
$+v(z,x_{i})\phi(w-z,x_{i}+x_{j})$ $+$ $v(-z,x_{j})\phi(w+z,x_{i}+x_{j})\,,$
and
$v(q_{j},x_{i})\phi(q_{i}-q_{j},x_{i}-x_{j})=v(q_{i},x_{i})\phi(q_{i}-q_{j},-x_{i}-x_{j})$ $+$ (4.59)
$+v(q_{i},-x_{j})\phi(q_{i}+q_{j},x_{i}+x_{j})$ $+$ $v(q_{j},x_{j})\phi(q_{i}+q_{j},x_{i}-x_{j})\,.$
[…]
Below we formulate some relations which are used for the proof of the Lax equation in the next Section.
Corollary 4.2
The following identities hold true:
$\begin{array}[]{c}\displaystyle{{R}^{z}_{ij}{F}^{z}_{jk}-{F}^{z}_{ij}{R}^{z}_{jk}={F}^{0}_{jk}R_{ik}^{z}-R_{ik}^{z}{F}^{0}_{ij}\,,}\end{array}$ (4.60)
$\begin{array}[]{c}\displaystyle{R^{z}_{ij}{\tilde{F}}^{z}_{jk}-{F}^{z}_{ij}{\tilde{R}}^{z}_{jk}={\tilde{F}}^{0}_{jk}{\tilde{R}}^{z}_{ik}-{\tilde{R}}^{z}_{ik}F^{0}_{ij}\,,}\end{array}$ (4.61)
$\begin{array}[]{c}\displaystyle{{\tilde{R}}^{z}_{ij}F^{-z}_{jk}+{\tilde{F}}^{z}_{ij}R^{-z}_{jk}=F^{0}_{jk}{\tilde{R}_{ik}}^{z}-{\tilde{R}_{ik}}^{z}{\tilde{F}}^{0}_{ij}\,,}\end{array}$ (4.62)
$\begin{array}[]{c}\displaystyle{{\tilde{R}}^{z}_{ij}{\tilde{F}}^{-z}_{jk}+{\tilde{F}}^{z}_{ij}{\tilde{R}}^{-z}_{jk}={\tilde{F}}^{0}_{jk}R_{ik}^{z}-R_{ik}^{z}{\tilde{F}}^{0}_{ij}\,,}\end{array}$ (4.63)
and
$\tilde{Y}_{j}^{0}{\tilde{R}}_{ij}^{z}-{\tilde{R}}_{ij}^{z}\tilde{Y}_{i}^{0}=-2{F}_{ij}^{z}\tilde{K}_{j}^{z}+2\tilde{K}_{i}^{z}{F}_{ij}^{-z}+{R}_{ij}^{z}\tilde{Y}_{j}^{z}+\tilde{Y}_{i}^{z}{R}_{ij}^{-z}\,,$ (4.64)
$\tilde{Y}_{j}^{0}R_{ij}^{z}-R_{ij}^{z}\tilde{Y}_{i}^{0}=2\tilde{F}_{ij}^{z}\tilde{K}_{j}^{-z}+2\tilde{K}_{i}^{z}\tilde{F}_{ij}^{-z}+\tilde{R}_{ij}^{z}\tilde{Y}_{j}^{-z}+\tilde{Y}_{i}^{z}\tilde{R}_{ij}^{-z}\,,$ (4.65)
where we use the short notations $R_{ij}^{\mu}:=R_{ij}^{\mu}(q_{i}-q_{j})$, $\tilde{R}_{ij}^{\mu}:=\tilde{R}_{ij}^{\mu}(q_{i}+q_{j})$, $\tilde{K}_{i}^{z}:=\tilde{K}_{i}^{z}(q_{i})$, $F_{ij}^{z}:=F_{ij}^{\mu}(q_{i}-q_{j})$ and $\tilde{F}_{ij}^{z}:=\tilde{F}_{ij}^{z}(q_{i}+q_{j})$, $\tilde{Y}_{i}^{z}:=\tilde{Y}_{i}^{z}(q_{i})$.
Proof. The identities (4.60), (4.61) are obtained from (4.42), (4.43) by differentiating with respect to $q_{j}$ and substitution $w:=z$. In the same manner the identities (4.62), (4.63) follow from (4.44), (4.45) by differentiating with respect to $q_{j}$ and substitutions $z:=0$ and $w:=z$.
In order to prove (4.64) we apply the operator $(\partial_{q_{i}}-\partial_{q_{j}})$ to (4.46):
$2F_{ij}^{w+z}\tilde{K}_{j}^{w}-R_{ij}^{w+z}\tilde{Y}_{j}^{w}=\tilde{Y}_{i}^{w}R_{ij}^{z-w}+2\tilde{K}_{i}^{w}F_{ij}^{z-w}+\tilde{R}_{ij}^{w-z}\tilde{Y}_{i}^{z}-\tilde{Y}_{j}^{-z}\tilde{R}_{ij}^{w+z}$ (4.66)
and put in $z:=0$ (4.66). By changing the indices $i\leftrightarrow j$ and the variables $q_{i}\leftrightarrow q_{j}$, $z\leftrightarrow-z$ in (4.46), and using the properties (4.37), (4.38), we obtain:
$\tilde{K}_{j}^{w}R_{ij}^{w+z}=R_{ij}^{-w+z}\tilde{K}_{i}^{w}+\tilde{R}^{w+z}_{ij}\tilde{K}_{j}^{-z}+\tilde{K}_{i}^{z}\tilde{R}_{ij}^{w-z}\,.$ (4.67)
Next, by applying the operator $(\partial_{q_{i}}+\partial_{q_{j}})$ to (4.67) one gets:
$\tilde{Y}_{j}^{w}R_{ij}^{w+z}=R_{ij}^{-w+z}\tilde{Y}_{i}^{w}+2\tilde{F}_{ij}^{w+z}\tilde{K}_{j}^{-z}+2\tilde{K}_{i}^{z}\tilde{F}_{ij}^{w-z}+\tilde{R}_{ij}^{w+z}\tilde{Y}_{j}^{-z}+\tilde{Y}_{i}^{z}\tilde{R}_{ij}^{w-z}\,.$ (4.68)
[…]
For the proof of the Lax equation in the next Section we also use the identity
$\begin{array}[]{c}\displaystyle{R^{z}_{ij}{\tilde{R}}^{z}_{ji}={\tilde{R}}^{z}_{ji}R^{z}_{ij}\,.}\end{array}$ (4.69)
5 $R$-matrix-valued generalization of Takasaki’s Lax pair
5.1 The Lax pair
[…]
where all entries ${\mathcal{L}}^{ab}(z)$ and ${\mathcal{M}}^{ab}(z)$, $a,b=1,2$ are as follows:
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{11}_{ij}(z)=\delta_{ij}p_{i}+g(1-\delta_{ij})R_{ij}^{z}(q_{ij})\,,}\end{array}$ (5.2)
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{12}_{ij}(z)=\delta_{ij}\tilde{K}^{z}_{i}(q_{i})+g(1-\delta_{ij})\tilde{R}^{z}_{ij}(q^{+}_{ij})\,,}\end{array}$ (5.3)
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{21}_{ij}(z)=-\delta_{ij}\tilde{K}^{-z}_{i}(q_{i})-g(1-\delta_{ij})\tilde{R}^{-z}_{ij}(q^{+}_{ij})\,,}\end{array}$ (5.4)
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{22}_{ij}(z)=-\delta_{ij}p_{i}-g(1-\delta_{ij})R^{-z}_{ij}(q_{ij})}\end{array}$ (5.5)
and
$\begin{array}[]{c}\displaystyle{{\mathcal{M}}^{11}_{ij}(z)=\delta_{ij}A_{i}+g(1-\delta_{ij})F^{z}_{ij}(q_{ij})\,,}\end{array}$ (5.6)
$\begin{array}[]{c}\displaystyle{{\mathcal{M}}^{12}_{ij}(z)=\frac{1}{2}\,\delta_{ij}\tilde{Y}_{i}^{z}(q_{i})+g(1-\delta_{ij})\tilde{F}^{z}_{ij}(q^{+}_{ij})\,,}\end{array}$ (5.7)
$\begin{array}[]{c}\displaystyle{{\mathcal{M}}^{21}_{ij}(z)=\frac{1}{2}\,\delta_{ij}\tilde{Y}^{-z}_{i}(q_{i})+g(1-\delta_{ij})\tilde{F}^{-z}_{ij}(q^{+}_{ij})\,,}\end{array}$ (5.8)
$\begin{array}[]{c}\displaystyle{{\mathcal{M}}^{22}_{ij}(z)=\delta_{ij}A_{i}+g(1-\delta_{ij})F^{-z}_{ij}(q_{ij})\,.}\end{array}$ (5.9)
Here
$\begin{array}[]{c}\displaystyle{A_{i}=-\frac{1}{2}\,\tilde{Y}_{i}^{0}(q_{i})-g\sum_{k:k\neq i}^{n}\Big{(}F^{0}_{ik}(q_{ik})+{\tilde{F}}^{0}_{ik}(q_{ik}^{+})\Big{)}}\end{array}$ (5.10)
and
$\begin{array}[]{c}\displaystyle{{\mathcal{H}}=g\sum_{k<l}^{n}\Big{(}F^{0}_{kl}(q_{kl})+{\tilde{F}}^{0}_{kl}(q_{kl}^{+})\Big{)}+\frac{1}{2}\sum\limits_{k=1}^{n}\tilde{Y}_{k}^{0}(q_{k})\,.}\end{array}$ (5.11)
In the next subsection we prove the following statement.
Theorem 5.1
The Lax pair (5.1)-(5.11) satisfies the Lax equation
˙$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),{\mathcal{M}}(z)]}\end{array}$ (5.12)
and provides the equations of motion for the classical Calogero-Inozemtsev model (3.1).
5.2 Proof of the Lax equation
Here we generalize the calculations from the Section 3.2.
The block 11.
The form of the equations in this case is similar to (3.26), but this time it is necessary to take into account the additional term ${\mathcal{H}}$ in the diagonal part of ${\mathcal{M}}$. Consequently, for $a=1,2$ we have
~$\begin{array}[]{c}\displaystyle{\tilde{{\mathcal{M}}}^{aa}_{ii}(z):={\mathcal{M}}^{aa}_{ii}(z)+{\mathcal{H}}=A_{i}+{\mathcal{H}}=\frac{1}{2}\sum\limits_{k,l\neq i}^{n}g\Big{(}F^{0}_{kl}(q_{kl})+{\tilde{F}}^{0}_{kl}(q_{kl}^{+})\Big{)}+\frac{1}{2}\sum\limits_{k\neq i}^{n}{\tilde{Y}}_{k}^{0}(q_{k})\,,}\end{array}$ (5.13)
where we used the properties
$F_{kl}^{0}=F_{lk}^{0}\qquad\text{and}\qquad\tilde{F}_{kl}^{0}=\tilde{F}_{lk}^{0}$ (5.14)
following from (4.37) and (4.38).
The diagonal part takes the form:
$\begin{array}[]{c}\displaystyle{\dot{{\mathcal{L}}^{11}_{ii}}={\mathcal{L}}^{11}_{ii}{\tilde{\mathcal{M}}}^{11}_{ii}-{\tilde{\mathcal{M}}}^{11}_{ii}{\mathcal{L}}^{11}_{ii}+{\mathcal{L}}^{12}_{ii}{\mathcal{M}}^{21}_{ii}-{\mathcal{M}}^{12}_{ii}{\mathcal{L}}^{21}_{ii}+}\\ \\ \displaystyle{+\sum\limits_{k:k\neq i}^{n}\Big{(}{\mathcal{L}}^{12}_{ik}{\mathcal{M}}^{21}_{ki}-{\mathcal{M}}^{12}_{ik}{\mathcal{L}}^{21}_{ki}+{\mathcal{L}}^{11}_{ik}{\mathcal{M}}^{11}_{ki}-{\mathcal{M}}^{11}_{ik}{\mathcal{L}}^{11}_{ki}\Big{)}\,.}\end{array}$ (5.15)
Plugging the elements of ${\mathcal{L}}$- and ${\mathcal{M}}$-operators into (5.15), we obtain the equation of motion (3.1):
$\begin{array}[]{c}\displaystyle{\dot{p_{i}}=p_{i}(A_{i}+{\mathcal{H}})-(A_{i}+{\mathcal{H}})p_{i}+\frac{1}{2}{\tilde{K}}^{z}_{i}{\tilde{Y}}^{-z}_{i}+\frac{1}{2}{\tilde{Y}}^{z}_{i}{\tilde{K}}^{-z}_{i}+}\\ \\ \displaystyle{+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}{\tilde{R}}^{z}_{ik}{\tilde{F}}^{-z}_{ki}+{\tilde{F}}^{z}_{ik}{\tilde{R}}^{-z}_{ki}+R^{z}_{ik}F^{z}_{ki}-F^{z}_{ik}R^{z}_{ki}\Big{)}=}\\ \\ \displaystyle{=\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(q_{i}+\omega_{a})+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}\wp^{\prime}(q^{+}_{ik})+\wp^{\prime}(q_{ik})\Big{)}\,,}\end{array}$ (5.16)
where we used (4.28), (4.26), (4.25).
For the non-diagonal part we need to show
$\begin{array}[]{c}\displaystyle{\dot{{\mathcal{L}}^{11}_{ij}}={\mathcal{L}}^{11}_{ii}{\mathcal{M}}^{11}_{ij}-{\mathcal{M}}^{11}_{ij}{\mathcal{L}}^{11}_{jj}+{\mathcal{L}}^{12}_{ij}{\mathcal{M}}^{21}_{jj}-{\mathcal{M}}^{12}_{ij}{\mathcal{L}}^{21}_{jj}+{\mathcal{L}}^{12}_{ii}{\mathcal{M}}^{21}_{ij}-{\mathcal{M}}^{12}_{ii}{\mathcal{L}}^{21}_{ij}+}\\ \\ \displaystyle{+{\mathcal{L}}^{11}_{ij}\tilde{{\mathcal{M}}}^{11}_{jj}-\tilde{{\mathcal{M}}}^{11}_{ii}{\mathcal{L}}^{11}_{ij}+\sum\limits_{k:k\neq i,j}^{n}\Big{(}{\mathcal{L}}^{12}_{ik}{\mathcal{M}}^{21}_{kj}-{\mathcal{M}}^{12}_{ik}{\mathcal{L}}^{21}_{kj}+{\mathcal{L}}^{11}_{ik}{\mathcal{M}}^{11}_{kj}-{\mathcal{M}}^{11}_{ik}{\mathcal{L}}^{11}_{kj}\Big{)}\,.}\end{array}$ (5.17)
Here we use the notation ~$\tilde{{\mathcal{M}}}_{ii}^{aa}$ from (5.13). The l.h.s. of (5.17) equals $g(\dot{q_{i}}-\dot{q_{j}})F^{z}_{ij}$. For the r.h.s. of (5.17) we have:
$\begin{array}[]{c}\displaystyle{g(p_{i}-p_{j})F^{z}_{ij}+\frac{g}{2}\,{\tilde{R}}^{z}_{ij}{\tilde{Y}}^{-z}_{j}+g{\tilde{F}}^{z}_{ij}{\tilde{K}}^{-z}_{j}+g{\tilde{K}}^{z}_{i}{\tilde{F}}^{-z}_{ij}+\frac{g}{2}\,{\tilde{Y}}^{z}_{i}{\tilde{R}}^{-z}_{ij}+}\\ \\ \displaystyle{+\frac{g}{2}R^{z}_{ij}\Big{(}\sum\limits_{k,l\neq j}^{n}g\Big{(}F^{0}_{kl}+{\tilde{F}}^{0}_{kl}\Big{)}+\sum\limits_{k\neq j}^{n}{\tilde{Y}}_{k}^{0}\,\Big{)}-\frac{g}{2}\Big{(}\sum\limits_{k,l\neq i}^{n}g\Big{(}F^{0}_{kl}+{\tilde{F}}^{0}_{kl}\Big{)}+\sum\limits_{k\neq i}^{n}{\tilde{Y}}_{k}^{0}\,\Big{)}R^{z}_{ij}+}\\ \\ \displaystyle{+g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}{\tilde{R}}^{z}_{ik}{\tilde{F}}^{-z}_{kj}+{\tilde{F}}^{z}_{ik}{\tilde{R}}^{-z}_{kj}+R^{z}_{ik}F^{z}_{kj}-F^{z}_{ik}R^{z}_{kj}\Big{)}=}\end{array}$ (5.18)
$\begin{array}[]{c}\displaystyle{=g(p_{i}-p_{j})F^{z}_{ij}+g\left(\frac{1}{2}\,{\tilde{R}}^{z}_{ij}{\tilde{Y}}^{-z}_{j}+{\tilde{F}}^{z}_{ij}{\tilde{K}}^{-z}_{j}+\frac{1}{2}\,{\tilde{Y}}^{z}_{i}{\tilde{R}}^{-z}_{ij}+{\tilde{K}}^{z}_{i}{\tilde{F}}^{-z}_{ij}+\frac{1}{2}R^{z}_{ij}{\tilde{Y}}_{i}^{0}-\frac{1}{2}{\tilde{Y}}_{j}^{0}R^{z}_{ij}\right)+}\\ \\ \displaystyle{\underline{+g^{2}R^{z}_{ij}\sum\limits_{k:k\neq i,j}^{n}\Big{(}F^{0}_{ik}+{\tilde{F}}^{0}_{ik}\Big{)}-g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}F^{0}_{kj}+{\tilde{F}}^{0}_{kj}\Big{)}R^{z}_{ij}+}}\\ \\ \displaystyle{\underline{+g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}{\tilde{R}}^{z}_{ik}{\tilde{F}}^{-z}_{kj}+{\tilde{F}}^{z}_{ik}{\tilde{R}}^{-z}_{kj}+R^{z}_{ik}F^{z}_{kj}-F^{z}_{ik}R^{z}_{kj}\Big{)}}=}\\ \\ \displaystyle{=g(p_{i}-p_{j})F^{z}_{ij}\,.}\end{array}$
In the first equality we used the commutativity properties (4.39), (4.40), (4.41) and their consequences $F_{kl}^{0}R_{ij}^{z}=R_{ij}^{z}F_{kl}^{0}$ and $\tilde{F}_{kl}^{0}R_{ij}^{z}=R_{ij}^{z}\tilde{F}_{kl}^{0}$ for distinct $i,j,k,l$, so that the corresponding terms with distinct indices vanish. Also, we used (5.14). The terms inside the brackets vanish due to (4.65). The underlined terms are cancelled out due to (4.62) and (4.60).
The block 12.
In the diagonal part we need to prove the following equation
$\begin{array}[]{c}\displaystyle{\dot{{\mathcal{L}}^{12}_{ii}}={\mathcal{L}}^{11}_{ii}{\mathcal{M}}^{12}_{ii}-{\mathcal{M}}^{12}_{ii}{\mathcal{L}}^{22}_{ii}+{\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}-\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}+}\\ \\ \displaystyle{+\sum\limits_{k:k\neq i}^{n}\Big{(}{\mathcal{L}}^{11}_{ik}{\mathcal{M}}^{12}_{ki}-{\mathcal{M}}^{11}_{ik}{\mathcal{L}}^{12}_{ki}+{\mathcal{L}}^{12}_{ik}{\mathcal{M}}^{22}_{ki}-{\mathcal{M}}^{12}_{ik}{\mathcal{L}}^{22}_{ki}\Big{)}\,.}\end{array}$ (5.19)
Here we have additional terms ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}-\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$, which are absent in (3.32). However, the equality ${\mathcal{L}}_{ii}^{12}\tilde{\mathcal{M}}_{ii}^{22}=\tilde{\mathcal{M}}_{ii}^{11}{\mathcal{L}}_{ii}^{12}$ holds true in the operator case as well. This follows from the explicit form of ~$\tilde{\mathcal{M}}_{ii}^{aa}$ given in (5.13) and (4.39), (4.40). Therefore, similarly to (3.33) we have:
$\begin{array}[]{c}\displaystyle{\dot{q_{i}}{\tilde{Y}}^{z}_{i}=p_{i}{\tilde{Y}}^{z}_{i}+g^{2}\sum\limits_{k:k\neq i}^{n}\Big{(}R^{z}_{ik}(q_{ik}){\tilde{F}}^{z}_{ki}(q^{+}_{ik})+{\tilde{F}}^{z}_{ik}(q^{+}_{ik})R^{-z}_{ki}(q_{ki})+}\\ \\ \displaystyle{+{\tilde{R}}^{z}_{ik}(q^{+}_{ik})F^{-z}_{ki}(q_{ki})-F^{z}_{ik}(q_{ik}){\tilde{R}}^{z}_{ki}(q^{+}_{ki})\Big{)}=p_{i}{\tilde{Y}}^{z}_{i}\,,}\end{array}$ (5.20)
where we used the commutativity [^3] A sufficient condition is that $\frac{\partial}{\partial q_{k}}\left(R_{ik}^{z}(q_{i}-q_{k})\tilde{R}^{z}_{ik}(q_{i}+q_{k})-\tilde{R}_{ik}^{z}(q_{i}+q_{k})R^{z}_{ik}(q_{i}-q_{k})\right)=0$. (4.69) and the skew-symmetry (4.38).
For the non-diagonal part we have the equation:
$\begin{array}[]{c}\displaystyle{\dot{{\mathcal{L}}^{12}_{ij}}={\mathcal{L}}^{11}_{ii}{\mathcal{M}}^{12}_{ij}-{\mathcal{M}}^{12}_{ij}{\mathcal{L}}^{22}_{jj}+{\mathcal{L}}^{11}_{ij}{\mathcal{M}}^{12}_{jj}-{\mathcal{M}}^{11}_{ij}{\mathcal{L}}^{12}_{jj}+{\mathcal{L}}^{12}_{ii}{\mathcal{M}}^{22}_{ij}-{\mathcal{M}}^{12}_{ii}{\mathcal{L}}^{22}_{ij}+{\mathcal{L}}^{12}_{ij}\tilde{\mathcal{M}}^{22}_{jj}-\tilde{{\mathcal{M}}}^{11}_{ii}{\mathcal{L}}^{12}_{ij}}\\ \\ \displaystyle{+\sum\limits_{k:k\neq i,j}^{n}\Big{(}{\mathcal{L}}^{11}_{ik}{\mathcal{M}}^{12}_{kj}-{\mathcal{M}}^{11}_{ik}{\mathcal{L}}^{12}_{kj}+{\mathcal{L}}^{12}_{ik}{\mathcal{M}}^{22}_{kj}-{\mathcal{M}}^{12}_{ik}{\mathcal{L}}^{22}_{kj}\Big{)}\,.}\end{array}$ (5.21)
Similarly to (5.18), the l.h.s. of (5.21) equals $g(\dot{q_{i}}+\dot{q_{j}}){\tilde{F}}^{z}_{ij}$. For the r.h.s. of (5.21) we have:
$\begin{array}[]{c}\displaystyle{g(p_{i}+p_{j}){\tilde{F}}^{z}_{ij}+\frac{g}{2}\,{R}^{z}_{ij}{\tilde{Y}}^{z}_{j}-g{F}^{z}_{ij}{\tilde{K}}^{z}_{j}+g{\tilde{K}}^{z}_{i}{F}^{-z}_{ij}+\frac{g}{2}\,{\tilde{Y}}^{z}_{i}{R}^{-z}_{ij}+}\\ \\ \displaystyle{+\frac{g}{2}{\tilde{R}}^{z}_{ij}\Big{(}\sum\limits_{k,l\neq j}^{n}g\Big{(}F^{0}_{kl}+{\tilde{F}}^{0}_{kl}\Big{)}+\sum\limits_{k\neq j}^{n}{\tilde{Y}}_{k}^{0}\,\Big{)}-\frac{g}{2}\Big{(}\sum\limits_{k,l\neq i}^{n}g\Big{(}F^{0}_{kl}+{\tilde{F}}^{0}_{kl}\Big{)}+\sum\limits_{k\neq i}^{n}{\tilde{Y}}_{k}^{0}\,\Big{)}{\tilde{R}}^{z}_{ij}+}\\ \\ \displaystyle{+g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}R^{z}_{ik}{\tilde{F}}^{z}_{kj}-F^{z}_{ik}{\tilde{R}}^{z}_{kj}+{\tilde{R}}^{z}_{ik}{F}^{-z}_{kj}+{\tilde{F}}^{z}_{ik}{R}^{-z}_{kj}\Big{)}=}\end{array}$ (5.22)
$\begin{array}[]{c}\displaystyle{=g(p_{i}+p_{j}){\tilde{F}}^{z}_{ij}+g\left(\frac{1}{2}\,{R}^{z}_{ij}{\tilde{Y}}^{z}_{j}-{F}^{z}_{ij}{\tilde{K}}^{z}_{j}+{\tilde{K}}^{z}_{i}{F}^{-z}_{ij}+\frac{1}{2}\,{\tilde{Y}}^{z}_{i}{R}^{-z}_{ij}+\frac{1}{2}{\tilde{R}}^{z}_{ij}{\tilde{Y}}_{i}^{0}-\frac{1}{2}{\tilde{Y}}_{j}^{0}{\tilde{R}}^{z}_{ij}\right)}\\ \\ \displaystyle{\underline{+g^{2}{\tilde{R}}^{z}_{ij}\sum\limits_{k:k\neq i,j}^{n}\Big{(}F^{0}_{ik}+{\tilde{F}}^{0}_{ik}\Big{)}-g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}F^{0}_{kj}+{\tilde{F}}^{0}_{kj}\Big{)}{\tilde{R}}^{z}_{ij}+}}\\ \\ \displaystyle{\underline{+g^{2}\sum\limits_{k:k\neq i,j}^{n}\Big{(}R^{z}_{ik}{\tilde{F}}^{z}_{kj}-F^{z}_{ik}{\tilde{R}}^{z}_{kj}+{\tilde{R}}^{z}_{ik}{F}^{-z}_{kj}+{\tilde{F}}^{z}_{ik}{R}^{-z}_{kj}\Big{)}}=}\\ \\ \displaystyle{=g(p_{i}+p_{j}){\tilde{F}}^{z}_{ij}\,.}\end{array}$
The terms inside the brackets vanish due to (4.64). The underlined terms are cancelled out due to (4.61), (4.62).
The block 21 and 22.
In the R-operator case, (anti)symmetry properties (3.12), (3.13) for ${\mathcal{L}}$, ${\mathcal{M}}$ remain valid:
${\mathcal{L}}^{11}(z)=-{\mathcal{L}}^{22}(-z),\qquad{\mathcal{L}}^{12}(z)=-{\mathcal{L}}^{21}(-z),$ (5.23)
~${\tilde{\mathcal{M}}}^{11}(z)={\tilde{\mathcal{M}}}^{22}(-z),\qquad{\mathcal{M}}^{12}(z)={\mathcal{M}}^{21}(-z).$ (5.24)
The equations
˙$\dot{{\mathcal{L}}}^{21}(z)={\mathcal{L}}^{21}(z){\tilde{\mathcal{M}}}^{11}(z)+{\mathcal{L}}^{22}(z){\mathcal{M}}^{21}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{11}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{21}(z),$ (5.25)
˙$\dot{{\mathcal{L}}}^{22}(z)={\mathcal{L}}^{21}(z){\mathcal{M}}^{12}(z)+{\mathcal{L}}^{22}(z){\tilde{\mathcal{M}}}^{22}(z)-{\mathcal{M}}^{21}(z){\mathcal{L}}^{12}(z)-{\tilde{\mathcal{M}}}^{22}(z){\mathcal{L}}^{22}(z)$ (5.26)
are correct due to (5.15), (5.17), (5.19), (5.21) and (5.23)-(5.24).
6 Representation via the Baxter’s 8-vertex $R$-matrix
The Baxter’s $R$-matrix has the form [2]:
$\begin{array}[]{c}\displaystyle{R_{12}^{\hbar}(z)=\frac{1}{2}\sum\limits_{k=0}^{3}\varphi_{k}\Big{(}z,\omega_{k}+\frac{\hbar}{2}\Big{)}\sigma_{4-k}\otimes\sigma_{4-k}\,,}\end{array}$ (6.1)
where
$\begin{array}[]{c}\displaystyle{\sigma_{0}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{1}\end{array}\right)\,,\quad\sigma_{1}=\left(\begin{array}[]{cc}{0}&{1}\\ {1}&{0}\end{array}\right)\,,\quad\sigma_{2}=\left(\begin{array}[]{cc}{0}&{-\imath}\\ {\imath}&{0}\end{array}\right)\,,\quad\sigma_{3}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{-1}\end{array}\right)}\end{array}$ (6.2)
are the Pauli matrices and $\sigma_{4}=\sigma_{0}=1_{2\times 2}$. Equivalently,
$\begin{array}[]{c}R_{12}^{\hbar}(z)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{cccc}\varphi_{00}+\varphi_{10}&0&0&\varphi_{01}-\varphi_{11}\\ 0&\varphi_{00}-\varphi_{10}&\varphi_{01}+\varphi_{11}&0\\ 0&\varphi_{01}+\varphi_{11}&\varphi_{00}-\varphi_{10}&0\\ \varphi_{01}-\varphi_{11}&0&0&\varphi_{00}+\varphi_{10}\end{array}\right)\,,\end{array}$ (6.3)
where
$\begin{array}[]{c}\displaystyle{\varphi_{00}=\phi\Big{(}z,\frac{\hbar}{2}\Big{)}=\varphi_{0}\Big{(}z,\omega_{0}+\frac{\hbar}{2}\Big{)}\,,\quad\varphi_{10}=\phi\Big{(}z,\frac{1+\hbar}{2}\Big{)}=\varphi_{1}\Big{(}z,\omega_{1}+\frac{\hbar}{2}\Big{)}\,,}\\ \\ \displaystyle{\varphi_{01}=e^{\pi\imath z}\phi\Big{(}z,\frac{\tau+\hbar}{2}\Big{)}=\varphi_{3}\Big{(}z,\omega_{3}+\frac{\hbar}{2}\Big{)}\,,\quad\varphi_{11}=e^{\pi\imath z}\phi\Big{(}z,\frac{1+\tau+\hbar}{2}\Big{)}=\varphi_{2}\Big{(}z,\omega_{2}+\frac{\hbar}{2}\Big{)}\,.}\end{array}$ (6.4)
Define also the $K$-matrix as
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{\hbar}(z)=\sum\limits_{k=0}^{3}\nu_{k}e^{-2\pi\imath\hbar\partial_{\tau}\omega_{k}}\varphi_{k}(z-\omega_{k},\hbar+\omega_{k})\sigma_{4-k}=}\\ \\ \displaystyle{=\sum\limits_{k=0}^{3}\nu_{k}e^{2\pi\imath(z+\hbar+\omega_{k})\partial_{\tau}\omega_{k}}\phi(z+\omega_{k},\hbar+\omega_{k})\sigma_{4-k}\,,}\end{array}$ (6.5)
where $\nu_{4}=\nu_{0}$ is assumed, and $\partial_{\tau}\omega_{k}$ is equal to either 00 (for $\omega_{0}$, $\omega_{1}$) or $1/2$ (for $\omega_{2}$, $\omega_{3}$). Up to simple redefinitions it coincides with the known $K$-matrix from [12, 17]. It satisfies the property (4.18) that can be deduced from (A.16) and the quasi-periodic behaviour (A.25)-(A.28). Notice also that ${\tilde{K}}^{\hbar}(z)={\tilde{K}}^{z}(\hbar)$, that is ${K}^{\hbar}(z)={\tilde{K}}^{\hbar}(z)$ in this description.
Proposition 6.1
Let $R^{\hbar}_{ij}(q)$ be the Baxter’s $R$-matrix (6.1). Then the $K$-matrix (6.5) satisfies the relations (4.46) and (4.65) with $R^{\hbar}_{ij}=R^{\hbar}_{ij}(q_{i}-q_{j})$ and ${\tilde{R}}^{\hbar}_{ij}=R^{\hbar}_{ij}(q_{i}+q_{j})$.
The proof is by straightforward calculation. One should use the set of identities, which are similar to those in (3.22).
In fact, all the properties or $R$-matrix and $K$-matrix necessary for the proof of the $R$-matrix valued Lax equation hold true.
Proposition 6.2
The Lax equation (1.4) with the Lax pair (5.1)-(5.11) defined through the Baxter’s elliptic $R$-matrix (6.1) and the $K$-matrix (6.5) is equivalent to the equations of motion (3.1).
Let us write down explicit form for the expression ${\mathcal{H}}$ (5.11):
$\begin{array}[]{c}\displaystyle{{\mathcal{H}}=g\sum_{k<l}^{n}\Big{(}F^{0}_{kl}(q_{k}-q_{l})+{\tilde{F}}^{0}_{kl}(q_{k}+q_{l})\Big{)}+\frac{1}{2}\sum\limits_{k=1}^{n}\tilde{Y}_{k}^{0}(q_{k})\in{\rm Mat}(2,\mathbb{C})^{\otimes n}\,,}\end{array}$ (6.6)
where
$\begin{array}[]{c}\displaystyle{F^{0}_{12}(x)={\tilde{F}}^{0}_{12}(x)=\partial_{x}R_{12}^{\hbar}(x)\Big{|}_{\hbar=0}=}\\ \\ \displaystyle{=-\frac{1}{2}\,E_{2}(x)\sigma_{0}\otimes\sigma_{0}+\frac{1}{2}\sum\limits_{a=1}^{3}\varphi_{a}(x,\omega_{a})\Big{(}E_{1}(x+\omega_{a})-E_{1}(x)-E_{1}(\omega_{a})\Big{)}\sigma_{4-a}\otimes\sigma_{4-a}}\end{array}$ (6.7)
and
$\begin{array}[]{c}\displaystyle{\tilde{Y}^{0}(x)=\partial_{x}K^{\hbar}(x)\Big{|}_{\hbar=0}=}\\ \\ \displaystyle{=-\nu_{0}E_{2}(x)\sigma_{0}-\sum\limits_{a=1}^{3}\nu_{a}\varphi_{a}(x-\omega_{a},\omega_{a})\Big{(}E_{1}(x+\omega_{a})-E_{1}(x)-E_{1}(\omega_{a})\Big{)}\sigma_{4-a}\,.}\end{array}$ (6.8)
7 Concluding remarks: XYZ long-range spin chain of ${\rm BC}_{n}$ type
As was explained in the end of Section 2 the classical $R$-matrix-valued Lax pair allows us to describe a quantum long-range spin chains. Consider some equilibrium position in the Calogero-Inozemtsev model
$\begin{array}[]{c}\displaystyle{p_{i}=0\,,\quad q_{i}=z_{i}\,,\ i=1,...,n}\end{array}$ (7.1)
defined by the equations ${\dot{p}}_{i}=0$, that is
${g^{2}\sum_{k:k\neq i}^{n}\Big{(}\wp^{\prime}(z_{i}-z_{k})+\wp^{\prime}(z_{i}+z_{k})\Big{)}+\frac{1}{2}\sum\limits_{a=0}^{3}\nu_{a}^{2}\wp^{\prime}(z_{i}+\omega_{a})=0\,,\quad i=1,...,n\,.}$ (7.2)
To the best of our knowledge, these equilibrium positions are unknown in the general case. At the same time, it is known that $z_{i}$ coincide with positions of zeros of solutions to the Heun equation (in the elliptic form) [8]. Consider the restriction of ${\mathcal{L}}(z)$ and ${\mathcal{M}}(z)$ (5.1)-(5.11) to the positions (7.1):
${{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{p_{j}=0,q_{j}=z_{j}}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=z_{j}}\,,\quad{{\mathcal{H}}^{0}}^{\prime}={\mathcal{H}}^{0}|_{q_{j}=z_{j}}\,.}$ (7.3)
Then we come to the quantum Lax equation
$\begin{array}[]{c}\displaystyle{[{{\mathcal{H}}^{0}}^{\prime},{\mathcal{L}}^{\prime}(z)]=[{\mathcal{L}}^{\prime}(z),{\mathcal{M}}^{\prime}(z)]\,,}\end{array}$ (7.4)
where the quantum Hamiltonian ${{\mathcal{H}}^{0}}^{\prime}$ is as follows:
$\begin{array}[]{c}\displaystyle{{{\mathcal{H}}^{0}}^{\prime}=g\sum_{k<l}^{n}\Big{(}F^{0}_{kl}(z_{k}-z_{l})+{\tilde{F}}^{0}_{kl}(z_{k}+z_{l})\Big{)}+\frac{1}{2}\sum\limits_{k=1}^{n}\tilde{Y}_{k}^{0}(z_{k})=}\\ \displaystyle{=\frac{g}{2}\sum_{k<l}^{n}\Big{[}-E_{2}(z_{k}-z_{l})\stackrel{{\scriptstyle k}}{{\sigma}}_{0}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{0}+}\\ \displaystyle{+\sum\limits_{a=1}^{3}\varphi_{a}(z_{k}-z_{l},\omega_{a})\Big{(}E_{1}(z_{k}-z_{l}+\omega_{a})-E_{1}(z_{k}-z_{l})-E_{1}(\omega_{a})\Big{)}\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\otimes\stackrel{{\scriptstyle l}}{{\sigma}}_{4-a}\Big{]}-}\\ \displaystyle{-\sum\limits_{k=1}^{n}\Big{[}\nu_{0}E_{2}(z_{k})\stackrel{{\scriptstyle k}}{{\sigma}}_{0}+\sum\limits_{a=1}^{3}\nu_{a}\varphi_{a}(z_{k}-\omega_{a},\omega_{a})\Big{(}E_{1}(z_{k}+\omega_{a})-E_{1}(z_{k})-E_{1}(\omega_{a})\Big{)}\stackrel{{\scriptstyle k}}{{\sigma}}_{4-a}\Big{]}\,.}\end{array}$ (7.5)
In fact, the existence of the Lax equation is not sufficient for integrability. However, we expect the model to be integrable, and we hope to prove this conjecture in future works. The properties of this model will be studied elsewhere.
8 Appendix: elliptic functions
Main definitions and properties.
We mainly deal with the elliptic Kronecker function [32]:
$\begin{array}[]{l}\displaystyle{\phi(z,u)=\frac{\vartheta^{\prime}(0)\vartheta(z+u)}{\vartheta(z)\vartheta(u)}=\phi(u,z)\,,\quad\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1\,,\quad\phi(-z,-u)=-\phi(z,u)\,,}\end{array}$ (A.1)
where $\vartheta(z)$ is the first Jacobi theta-function:
$\begin{array}[]{c}\displaystyle{\vartheta(z)=\vartheta(z,\tau)\equiv-\theta{\left[\begin{array}[]{c}1/2\\ 1/2\end{array}\right]}(z|\,\tau)\,,}\end{array}$ (A.2)
$\begin{array}[]{c}\displaystyle{\theta{\left[\begin{array}[]{c}a\\ b\end{array}\right]}(z|\,\tau)=\sum_{j\in{\zeta}}\exp\left(2\pi\imath(j+a)^{2}\frac{\tau}{2}+2\pi\imath(j+a)(z+b)\right)\,,\quad{\rm Im}(\tau)>0\,.}\end{array}$ (A.3)
Here $\tau$ is the moduli of elliptic curve $\mathbb{C}/(\mathbb{Z}+\tau\mathbb{Z})$, ${\rm Im}(\tau)>0$. It is sometimes useful to use the Jacobi theta-functions:
$\begin{array}[]{c}\displaystyle{\theta_{1}(u|\tau)=\vartheta(u,\tau)=-i\sum_{k\in{\zeta}}(-1)^{k}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},\qquad\theta_{2}(u|\tau)=\sum_{k\in{\zeta}}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},}\\ \\ \displaystyle{\theta_{3}(u|\tau)=\sum_{k\in{\zeta}}q^{k^{2}}e^{2\pi iku},\qquad\theta_{4}(u|\tau)=\sum_{k\in{\zeta}}(-1)^{k}q^{k^{2}}e^{2\pi iku},}\end{array}$ (A.4)
where $q=e^{\pi i\tau}$ and
$\begin{array}[]{c}\theta_{2}(u|\tau)=\theta_{1}(u+\frac{1}{2}|\tau),\!\quad\!\!\theta_{3}(u|\tau)=q^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau+1}{2}|\tau),\!\quad\!\!\theta_{4}(u|\tau)=-iq^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau}{2}|\tau).\end{array}$ (A.5)
The infinite product representation is
$\begin{array}[]{l}\displaystyle{\theta_{1}(u|\tau)=2q^{\frac{1}{4}}\sin\pi u\prod_{n\geq 1}(1-q^{2n})(1-q^{2n}e^{2\pi iu})(1-q^{2n}e^{-2\pi iu}),}\\ \\ \displaystyle{\theta_{2}(u|\tau)=2q^{\frac{1}{4}}\cos\pi u\prod_{n\geq 1}(1-q^{2n})(1+q^{2n}e^{2\pi iu})(1+q^{2n}e^{-2\pi iu}),}\\ \\ \displaystyle{\theta_{3}(u|\tau)=\prod_{n\geq 1}(1-q^{2n})(1+q^{2n-1}e^{2\pi iu})(1+q^{2n-1}e^{-2\pi iu}),}\\ \\ \displaystyle{\theta_{4}(u|\tau)=\prod_{n\geq 1}(1-q^{2n})(1-q^{2n-1}e^{2\pi iu})(1-q^{2n-1}e^{-2\pi iu}).}\end{array}$ (A.6)
The partial derivative $f(z,u)=\partial_{u}\varphi(z,u)$ equals
$\begin{array}[]{c}\displaystyle{f(z,u)=\phi(z,u)(E_{1}(z+u)-E_{1}(u)),\qquad f(-z,-u)=f(z,u)\,,}\end{array}$ (A.7)
where $E_{1}(z)$ the first Eisenstein function. The first and the second Eisenstein functions are defined as follows:
$\begin{array}[]{c}\displaystyle{E_{1}(z)=\frac{\vartheta^{\prime}(z)}{\vartheta(z)}=\zeta(z)+\frac{z}{3}\frac{\vartheta^{\prime\prime\prime}(0)}{\vartheta^{\prime}(0)}\,,\quad E_{2}(z)=-\partial_{z}E_{1}(z)=\wp(z)-\frac{\vartheta^{\prime\prime\prime}(0)}{3\vartheta^{\prime}(0)}\,,}\end{array}$ (A.8)
$\begin{array}[]{c}\displaystyle{E_{1}(-z)=-E_{1}(z)\,,\quad E_{2}(-z)=E_{2}(z)\,,}\end{array}$ (A.9)
where $\wp(z)$ and $\zeta(z)$ are the Weierstrass functions. Due to the following behavior of $\phi(z,q)$ near $z=0$
$\begin{array}[]{l}\displaystyle{\phi(z,q)=z^{-1}+E_{1}(q)+z\,(E_{1}^{2}(q)-\wp(q))/2+O(z^{2})\,.}\end{array}$ (A.10)
we also have
$\begin{array}[]{l}\displaystyle{f(0,q)=-E_{2}(q)\,.}\end{array}$ (A.11)
We use the following addition formulae:
$\begin{array}[]{c}\displaystyle{\phi(z_{1},u_{1})\phi(z_{2},u_{2})=\phi(z_{1},u_{1}+u_{2})\phi(z_{2}-z_{1},u_{2})+\phi(z_{2},u_{1}+u_{2})\phi(z_{1}-z_{2},u_{1})\,,}\end{array}$ (A.12)
and by applying $\partial_{u_{2}}-\partial_{u_{1}}$ we get
$\begin{array}[]{c}\displaystyle{\phi(z_{1},u_{1})f(z_{2},u_{2})-f(z_{1},u_{1})\phi(z_{2},u_{2})=}\\ \\ \displaystyle{=\phi(z_{1},u_{1}+u_{2})f(z_{2}-z_{1},u_{2})-\phi(z_{2},u_{1}+u_{2})f(z_{1}-z_{2},u_{1})\,.}\end{array}$ (A.13)
Also,
$\begin{array}[]{c}\displaystyle{\phi(z,u_{1})\phi(z,u_{2})=\phi(z,u_{1}+u_{2})\Big{(}E_{1}(z)+E_{1}(u_{1})+E_{1}(u_{2})-E_{1}(z+u_{1}+u_{2})\Big{)}\,,}\end{array}$ (A.14)
$\begin{array}[]{c}\displaystyle{\phi(z,u_{1})f(z,u_{2})-\phi(z,u_{2})f(z,u_{1})=\phi(z,u_{1}+u_{2})\Big{(}\wp(u_{1})-\wp(u_{2})\Big{)}=}\\ \\ \displaystyle{=\phi(z,u_{1}+u_{2})\Big{(}E_{2}(u_{1})-E_{2}(u_{2})\Big{)}=\phi(z,u_{1}+u_{2})\Big{(}f(0,u_{2})-f(0,u_{1})\Big{)}\,.}\end{array}$ (A.15)
$\begin{array}[]{c}\displaystyle{\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)=E_{2}(z)-E_{2}(u)\,,}\end{array}$ (A.16)
$\begin{array}[]{c}\displaystyle{\phi(z,u)f(z,-u)-\phi(z,-u)f(z,u)=\wp^{\prime}(u)\,.}\end{array}$ (A.17)
Define
$\begin{array}[]{c}\displaystyle{\omega_{0}=0\,,\quad\omega_{1}=\frac{1}{2}\,,\quad\omega_{2}=\frac{1+\tau}{2}\,,\quad\omega_{3}=\frac{\tau}{2}}\end{array}$ (A.18)
and
$\begin{array}[]{c}\displaystyle{\varphi_{0}(z,\hbar)=\phi(z,\hbar)\,,\qquad\varphi_{1}(z,\hbar+\omega_{1})=\phi(z,\omega_{1}+\hbar)\,,}\\ \\ \displaystyle{\varphi_{2}(z,\hbar+\omega_{2})=e^{\pi\imath z}\phi(z,\omega_{2}+\hbar)\,,\qquad\varphi_{3}(z,\hbar+\omega_{3})=e^{\pi\imath z}\phi(z,\omega_{3}+\hbar)\,.}\end{array}$ (A.19)
Then
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,\hbar+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{k+1}(z+\hbar)}{\theta_{1}(z)\theta_{k+1}(\hbar)}\,,\quad k=0,1,2,3}\end{array}$ (A.20)
The set of introduced functions is transformed under the action of matrix $I$ (3.18)-(A.22) as
$\begin{array}[]{c}\left(\begin{array}[]{c}\varphi_{0}(2z,u+\omega_{0})\\ \varphi_{1}(2z,u+\omega_{1})\\ \varphi_{2}(2z,u+\omega_{2})\\ \varphi_{3}(2z,u+\omega_{3})\end{array}\right)=\displaystyle{\frac{1}{2}}\left(\begin{array}[]{cccc}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right)\left(\begin{array}[]{c}\varphi_{0}(2u,z+\omega_{0})\\ \varphi_{1}(2u,z+\omega_{1})\\ \varphi_{2}(2u,z+\omega_{2})\\ \varphi_{3}(2u,z+\omega_{3})\end{array}\right)\,.\end{array}$ (A.21)
This transformation matrix can be written as
$\begin{array}[]{c}\displaystyle{I_{km}=\frac{1}{2}\exp\Big{(}4\pi\imath\Big{(}\omega_{m-1}\partial_{\tau}\omega_{k-1}-\omega_{k-1}\partial_{\tau}\omega_{m-1}\Big{)}\Big{)}\,,\quad k,m=1,...,4\,.}\end{array}$ (A.22)
It has the property
$\begin{array}[]{c}\displaystyle{I^{-1}=I\,.}\end{array}$ (A.23)
We also use the widely known identity:
$\begin{array}[]{c}\displaystyle{\sum\limits_{a=0}^{3}\wp(z+\omega_{a})=4\wp(2z)\,.}\end{array}$ (A.24)
The quasi-periodic properties are as follows:
$\begin{array}[]{l}\displaystyle{\phi(z+1,u)=\phi(z,u)\,,\qquad\phi(z+\tau,u)=e^{-2\pi\imath u}\phi(z,u)}\end{array}$ (A.25)
or
$\begin{array}[]{l}\displaystyle{\phi(z\pm 2\omega_{a},u)=e^{\mp 4\pi\imath\partial_{\tau}\omega_{a}u}\phi(z,u)\,.}\end{array}$ (A.26)
For the theta function (2.7) we have
$\begin{array}[]{l}\displaystyle{\vartheta(z+2\omega_{a})=-e^{-4\pi\imath(z+\omega_{a})\partial_{\tau}\omega_{a}}\vartheta(z)}\end{array}$ (A.27)
or
$\begin{array}[]{l}\displaystyle{\vartheta(z+\omega_{a})=-e^{-4\pi\imath z\partial_{\tau}\omega_{a}}\vartheta(z-\omega_{a})\,.}\end{array}$ (A.28)
This provides
$\begin{array}[]{l}\displaystyle{E_{2}(z+2\omega_{a})=E_{2}(z)\,,\qquad E_{1}(z+2\omega_{a})=E_{1}(z)-4\pi\imath\partial_{\tau}\omega_{a}}\end{array}$ (A.29)
and
$\begin{array}[]{l}\displaystyle{E_{1}(\omega_{a})=-2\pi\imath\partial_{\tau}\omega_{a}\,,\quad a\neq 0\,.}\end{array}$ (A.30)
Function $v(z,u)$ and related identities.
Here we briefly discuss identities related to the function $v(z,u)$ (3.15). It follows from (3.19) and (3.17) that
$v(z,u)v(-z,u)=\sum\limits_{a=0}^{3}\Big{(}\nu_{a}^{2}\wp(u+\omega_{a})-\bar{\nu}_{a}^{2}\wp(z+\omega_{a})\Big{)}=-v(z,u)v(z,-u)\,.$ (A.31)
Let us prove the important identities (3.22) and (3.23). The identity (3.22) consists of four relations proportional to $\nu_{0},...,\nu_{3}$. Consider the one proportional to $\nu_{0}$:
$\begin{array}[]{c}\displaystyle{\phi(2x,u)\phi(x+y,w-u)+\phi(2x,w)\phi(x-y,u-w)+\phi(2y,-u)\phi(x+y,u+w)=\,.}\\ \\ \displaystyle{=\phi(2y,w)\phi(x-y,u+w)\,.}\end{array}$ (A.32)
Using (A.12) we have
$\begin{array}[]{c}\displaystyle{\phi(x+y,w)\phi(x-y,u)=\phi(2x,u)\phi(x+y,w-u)-\phi(2x,w)\phi(y-x,w-u)\,.}\end{array}$ (A.33)
On the other hand for the same function we also have
$\begin{array}[]{c}\displaystyle{\phi(x+y,w)\phi(x-y,u)=\phi(2y,w)\phi(x-y,u+w)-\phi(2y,-u)\phi(x+y,u+w)\,.}\end{array}$ (A.34)
By comparing left hand sides of (A.33) and (A.34) one gets (A.32). For the rest of the components of (3.22) the proof is similar.
In order to prove (3.23) let us compute the partial derivative $\partial_{u}+\partial_{w}$ in (3.22):
$\begin{array}[]{c}v^{\prime}(x,u)\phi(x+y,w-u)+v^{\prime}(x,w)\phi(x-y,u-w)-\\ \\ -v^{\prime}(y,-u)\phi(x+y,u+w)+2v(y,-u)f(x+y,u+w)=\\ \\ =v^{\prime}(y,w)\phi(x-y,u+w)+2v(y,w)f(x-y,u+w)\end{array}$ (A.35)
and then we put $x=0$ and $y:=-z$:
$\begin{array}[]{c}v^{\prime}(0,u)\phi(-z,w-u)+v^{\prime}(0,w)\phi(z,u-w)-v^{\prime}(-z,-u)\phi(-z,u+w)+\\ \\ +2v(-z,-u)f(-z,u+w)=v^{\prime}(-z,w)\phi(z,u+w)+2v(-z,w)f(z,u+w)\,.\end{array}$ (A.36)
Finally, one should use an obvious antisymmetry property
$\phi(z,u)=-\phi(-z,-u),\qquad v(-z,-u)=-v(z,u).$ (A.37)
to get (3.23).
Acknowledgments
We are grateful to O. Chalykh, An. Kirillov, A. Liashyk, V. Prokofev and I. Sechin for useful discussions.
This work was supported by the Russian Science Foundation under grant no. 25-11-00081,
https://rscf.ru/en/project/25-11-00081/ and performed at Steklov Mathematical Institute of Russian Academy of Sciences.